### v1

# Lower bounds for the eigenvalues of the Hodge Laplacian on certain non-convex domains

Abstract.
We establish geometric lower bounds for the smallest positive eigenvalue of the Hodge Laplacian in the class of non-convex domains given by Euclidean annular regions with a convex outer boundary and a spherical inner boundary. These bounds are then extended to convex domains with multiple holes, where we derive lower bounds for certain higher order exact eigenvalues, and under additional geometric assumptions, also for the smallest positive eigenvalue. For $1$-forms on compact manifolds with boundary, we provide a general lower bound on the smallest exact eigenvalue – corresponding to the first positive Neumann eigenvalue – which, in certain respects, is better than the classical Cheeger inequality. Furthermore, we emphasise the necessity of the contact radius in the lower bounds of the main results. Our proofs employ local-to-global arguments via an explicit isomorphism between Cěch cohomology and de Rham cohomology to obtain Poincaré-type inequalities with explicit geometric dependence, and utilise certain generalised versions of the Cheeger–McGowan gluing lemma.
Key words and phrases:
Hodge Laplacian, eigenvalue bounds, differential forms, annular domains, Cěch cohomology, gluing lemma
2020 Mathematics Subject Classification:
58J50, 58A10, 58C40, 35P15
1. Introduction
The Hodge Laplacian on a Riemannian $n$-manifold $(M^{n},g)$, that acts on the space of smooth differential $p$-forms on $M$, for $p\in\{0,\dots,n\}$, is defined to be
$\Delta^{(p)}\coloneqq d\delta+\delta d,$
where $d$ denotes the exterior derivative and $\delta$ the codifferential. It is a generalisation of the Laplace-Beltrami operator acting on smooth functions on $M$ (i.e., 0-forms), and is well known for its interplay with geometry, analysis, and topology, besides its significance in modern mathematical physics. The spectral geometry of the Hodge Laplacian for $p>0$ is often more sophisticated and interesting compared to the Laplacian on functions, thanks to a greater interaction with the underlying topology arising from Hodge theory. For instance, the analogue of the classical Cheeger inequality which gives a lower bound for the spectrum in the case of functions, is not known in general for $p$-forms. See also the recent work [BC22] for a Cheeger-like lower bound on the smallest coexact eigenvalue for $1$-forms, which depends on certain curvature bounds.
Eigenvalue bounds for the Hodge Laplacian have been extensively studied in the literature. See [GM75, Dod82, CT97, Gue04a, Tak05, Sav09, RS11, Sav14, Kwo16, CHMN18, CS19, ECGH21] for some interesting lower bounds based on geometric and topological assumptions such as restrictions on the curvature, cohomology, codimension of isometric immersions, properties of certain ‘good covers’, and, in the case of manifolds with boundary, additional assumptions such as the star convexity of the domain or the $p$-convexity of the boundary. These type of bounds often employ some versions of the Bochner-Weitzenböck identity or the Reilly formula for differential forms. On the other hand, articles such as [CC90, PS01, Tak03, Lot04, Jam11, AT24] describe the construction of manifolds with small eigenvalues, the prescription of the $p$-spectrum under various restrictions, and the characterisation of manifolds with small eigenvalues.
In this article, we give a lower bound on the smallest positive eigenvalue of the Hodge Laplacian on $p$-forms with absolute boundary conditions, for a Euclidean annular domain whose outer boundary is convex and the inner boundary is a round sphere, and study its extension to a family of perforated domains with convex outer boundaries.
For convex Euclidean domains $\mathcal{D}\subset\mathbb{R}^{n}$, it was shown by the second named author [Gue04b] that the smallest positive absolute eigenvalue is bounded below by
$\mu^{(p)}_{1}(\mathcal{D})\geq\frac{C_{1}}{\operatorname{\mathrm{Diam}}(\mathcal{D})^{2}},\qquad\forall\,p\in\{0,\dots,n\},$
where $C_{1}$ is an explicit constant depending only on the dimension $n$ and the degree $p$ of the differential forms. A more refined version of these bounds were obtained by Savo [Sav11], also for convex Euclidean domains:
$\frac{C_{2}}{D_{p}(\mathcal{D})^{2}}\leq\mu^{(p)}_{1}(\mathcal{D})\leq\frac{C_{3}}{D_{p}(\mathcal{D})^{2}},\qquad\forall\,p\in\{1,\dots,n\},$
where $C_{2}$ and $C_{3}$ are explicit constants depending only on $n$ and $p$; $D_{p}(\mathcal{D})$ is the length of the $p^{\text{th}}$ longest principal axis of the largest ellipsoid that can be inscribed in $\mathcal{D}$, known as the John ellipsoid. For example, $D_{1}(\mathcal{D})$ corresponds to the diameter, and $D_{n}(\mathcal{D})$ corresponds to the inradius of $\mathcal{D}$. These bounds were also generalised for sufficiently small geodesically convex domains in manifolds with non-negative curvature operator [KK20].
Furthermore, for $p\in\{0,\dots,n-2\}$, Guerini [Gue04a, Theorem 2.1] constructed non-convex domains diffeomorphic to balls, with bounded diameters, which have an arbitrarily small positive eigenvalue for $p$-forms. The convexity assumption in [Gue04b] is a means to avoid the creation of non-trivial cohomology through collapsing as done in [Gue04a] by gluing a thick and a thin part that generalise the classical Cheeger dumbbell.
The lower bounds of this article for Euclidean annular regions depend on the distance between the inner hole and the outer boundary, which we refer to as the contact radius $R_{c}$.
Figure 1.
Notation. Throughout this article, $K$ denotes a generic constant depending only on the dimension $n$, possibly different in different expressions. We use $K_{i}$ to denote specific constants, also depending only on the dimension $n$.
We now state the main results.
Theorem 1.1.
Let $\mathfrak{D}\subset\mathbb{R}^{n}$ be a Euclidean annular domain, given by a smooth, bounded convex domain minus a hole. Then the following lower bounds hold for the smallest exact eigenvalue of the Hodge Laplacian on $p$-forms.
  1. (1)
    For $n\geq 3$ and $p\in\{2,\dots,n-1\}$,
    $\mu^{{(p)}}_{{1,\operatorname{\mathrm{ex}}}}(\mathfrak{D})\geq K\frac{1}{D^{2}}\left(\frac{R_{c}}{D}\right)^{2n^{2}-n-2}\min\left\{1,\left(\frac{R_{c}}{R_{h}}\right)^{3(n-1)(n+1)}\right\}.$
  2. (2)
    For $n\geq 2$ and $p=1$,
    $\mu^{{(1)}}_{{1,\operatorname{\mathrm{ex}}}}(\mathfrak{D})\geq K\frac{1}{D^{2}}\left(\frac{R_{c}}{D}\right)^{n-2}\min\left\{1,\left(\frac{R_{c}}{D}\right)^{3n+2}\left(\frac{R_{c}}{R_{h}}\right)^{7n-5}\right\}.$
Here, $D$ and $R_{c}$ denote the diameter and the contact radius of $\mathfrak{D}$, respectively, and $R_{h}$ the radius of the hole.
Remark 1.2.
The above lower bounds on $\mu^{{(p)}}_{{1,\operatorname{\mathrm{ex}}}}$ for each $p\in\{1,\dots,n-1\}$ will suffice to deduce lower bounds for the smallest positive absolute eigenvalue for all $p\in\{0,\dots,n\}$: For $p=0$, the smallest positive Neumann eigenvalue $\mu_{1}=\mu^{{(1)}}_{{1,\operatorname{\mathrm{ex}}}}$; for $p\in\{1,\dots,n-2\}$, we have
$\mu^{(p)}_{1}=\min\{\mu^{{(p)}}_{{1,\operatorname{\mathrm{ex}}}},\mu^{(p)}_{1,\operatorname{\mathrm{co}}}\}=\min\{\mu^{{(p)}}_{{1,\operatorname{\mathrm{ex}}}},\mu^{{(p+1)}}_{{1,\operatorname{\mathrm{ex}}}}\}$
in view of the correspondence between the coexact $p$-spectrum and the exact $(p+1)$-spectrum. Moreover, we have from the duality between the absolute and the relative spectra that
$\mu^{(n)}_{1}=\lambda_{1},$
where $\lambda_{1}$ is the first Dirichlet eigenvalue for functions, and
$\mu^{(n-1)}_{1}=\min\{\mu^{{(n-1)}}_{{1,\operatorname{\mathrm{ex}}}},\mu^{(n-1)}_{1,\operatorname{\mathrm{co}}}\}=\min\{\mu^{{(n-1)}}_{{1,\operatorname{\mathrm{ex}}}},\mu^{{(n)}}_{{1,\operatorname{\mathrm{ex}}}}\}=\min\{\mu^{{(n-1)}}_{{1,\operatorname{\mathrm{ex}}}},\mu^{(n)}_{1}\}=\min\{\mu^{{(n-1)}}_{{1,\operatorname{\mathrm{ex}}}},\lambda_{1}\},$
which covers the lower bounds on $\mu^{(n-1)}_{1}$ and $\mu^{(n)}_{1}$ using the Faber-Krahn inequality
$\lambda_{1}(\mathfrak{D})\geq\frac{K}{\operatorname{\mathrm{Diam}}(\mathfrak{D})^{2}}.$
To prove Theorem 1.1, we adapt the methods of Chanillo and Treves [CT97] – employing tools from Cěch–de Rham theory – in the scenario where the inner hole is sufficiently large compared to the contact radius; while for the case where the hole is much smaller than the contact radius, we use a combination of the bounds from the former case and the spectral convergence of Anne and Colbois [AC93], with the help of a useful gluing lemma by McGowan [McG93]. See Section 3.
We also prove the following theorem, which gives a better lower bound for the smallest positive Neumann eigenvalue than the one deduced from the Cheeger inequality. This theorem would be of use in deriving the lower bounds of Theorem 1.1 for the case of functions and $1$-forms.
Theorem 1.3.
Let $(M^{n},g)$ be a compact, connected, smooth Riemannian manifold with boundary ($n\geq 2$) and $\{U_{1},U_{2}\}$ be an open cover with $U_{1}$ and $U_{2}$ (each of which is connected) having smooth boundaries. Then the first positive Neumann eigenvalue satisfies
$\mu_{1}(M)\geq\frac{1}{32}\frac{\operatorname{\mathrm{Vol}}(U_{1}\cap U_{2})}{\operatorname{\mathrm{Vol}}(M)}\min\left\{\mu_{1}(U_{1}),\mu_{1}(U_{2})\right\}.$
Remark 1.4.
The theorem also holds for closed manifolds, with $\mu_{1}(M)$ on the LHS replaced by the first positive eigenvalue of the Laplace-Beltrami operator on $M$.
The above inequality is geometrically optimal in the following sense: It is known from [Ann95] that the first positive Neumann eigenvalue on a Euclidean Cheeger dumbbell with a cylinder of radius $\epsilon$} tends to 0 at a rate of ${\epsilon}^{n-1}$, which is same as the rate we get by applying our inequality with the open sets $U_{1}$ and $U_{2}$ each chosen to be the union of a ball and a cylinder of radius $\epsilon$}, because the ratio $\frac{\operatorname{\mathrm{Vol}}(U_{1}\cap U_{2})}{\operatorname{\mathrm{Vol}}(M)}$ too goes as ${\epsilon}^{n-1}$ and the eigenvalues $\mu_{1}$ of $U_{1}$ and $U_{2}$ are known to be bounded below independent of $\epsilon$}, from [Gue04a]. On the other hand, recall that the Cheeger inequality gives a rate of ${\epsilon}^{2(n-1)}$.
Figure 2.
Note, however, that the second positive Neumann eigenvalue of a manifold satisfies the lower bound [GHL04, Corollary 4.65]
$\mu_{2}(M)\geq\frac{1}{2}\min\left\{{\mu_{1}}(U_{1}),{\mu_{1}}(U_{2})\right\},$
and one would require a choice of $U_{1}$ and $U_{2}$ with small first positive eigenvalues in order to construct a manifold with small second positive eigenvalue.
Theorem 1.3 is proved in the latter half of Section 3, using only the variational characterisation for the Neumann eigenvalue and careful rearrangements of the terms involved.
Next, we show that the dependence of the lower bounds of Theorem 1.1 on the contact radius is necessary.
Theorem 1.5.
Let $n\geq 2$ and $p\in\{1,\dots,n-1\}$. There exists a family $(\mathcal{A}_{\epsilon}^{p})_{{\epsilon}\in(0,\frac{1}{2})}$ of Euclidean annular $C^{1}$-domains with convex outer boundary, such that $\lim_{{\epsilon}\to 0}\mu^{{(p)}}_{{1,\operatorname{\mathrm{ex}}}}(\mathcal{A}_{\epsilon}^{p})=0$, with the diameter of the domain and the radius of the hole uniformly bounded, and the contact radius tending to zero. It is also implied that the result holds for the smallest positive Neumann eigenvalue for functions, since $\mu_{1}=\mu^{{(1)}}_{{1,\operatorname{\mathrm{ex}}}}$.
The main idea of Theorem 1.5 involves the construction of an appropriate family of domains that tend to a limiting domain whose $p^{\text{th}}$ cohomology has a higher dimension. Then, using the harmonic form associated with the extra dimension of the cohomology on the limiting domain, we construct suitable test forms on these domains, to be used in the variational characterisation for the eigenvalues.
We now extend the lower bounds to the exact eigenvalues for convex Euclidean domains with multiple holes. We consider smooth, bounded domains $\mathfrak{D}\subset\mathbb{R}^{n}$ obtained by removing $\mathfrak{h}$} many non-overlapping balls (away from the boundary), of radius $R_{h}$ each, from a convex domain. Let $D$ denote the diameter of $\mathfrak{D}$, and define the contact radius of such domains with multiple holes to be
$\hat{R}_{c}(\mathfrak{D})\coloneqq\min\left\{r_{c},\frac{d_{h}}{2}\right\},$
where $r_{c}$ is the distance between the outer boundary and a hole closest to it, and $d_{h}$ is the distance between a pair of holes closest to each other.
Theorem 1.6.
Let $\mathfrak{D}\subset\mathbb{R}^{n}$ be a smooth, bounded domain as above, with convex outer boundary and $\mathfrak{h}$} many holes of radius $R_{h}$ each. Suppose that the Voronoi partition of $\mathfrak{D}$ with respect to the centers of the holes, up to reordering, is such that the union of the first $\ell$ pieces has a convex outer boundary, for every $\ell\in\{1,\dots,{\mathfrak{h}}\}$. Then we have the following lower bounds for the first exact eigenavalue.
  1. (i)
    For $n\geq 3$ and $p\in\{2,\dots,n-1\}$,
    $\mu^{{(p)}}_{{1,\operatorname{\mathrm{ex}}}}(\mathfrak{D})\geq K\frac{1}{D^{2}}\left(\frac{\hat{R}_{c}}{D}\right)^{2n^{2}-n-4+2{\mathfrak{h}}}\min\left\{1,\left(\frac{\hat{R}_{c}}{R_{h}}\right)^{3(n-1)(n+1)}\right\}.$
  2. (ii)
    For $n\geq 2$ and $p=1$,
    $\mu^{{(1)}}_{{1,\operatorname{\mathrm{ex}}}}(\mathfrak{D})\geq K\frac{1}{D^{2}}\left(\frac{\hat{R}_{c}R_{h}^{n-1}}{D^{n}}\right)^{{\mathfrak{h}}-1}\left(\frac{\hat{R}_{c}}{D}\right)^{n-2}\min\left\{1,\left(\frac{\hat{R}_{c}}{D}\right)^{3n+2}\left(\frac{\hat{R}_{c}}{R_{h}}\right)^{7n-5}\right\}.$
Although we use the Voronoi partition in the above theorem so as to have $\hat{R}_{c}$ defined independent of the partition, with an optimal contribution to the bounds, the lower bounds also hold if we consider any partition of $\mathfrak{D}$ into $\mathfrak{h}$} many pieces, each of which is a convex domain with a hole in the interior, and modify the definition of $\hat{R}_{c}$ appropriately. Further, similar lower bounds hold for a larger class of domains than those satisfying the above version of the geometric assumption. See Remark 5.4.
Remark 1.7.
We have stated Theorem 1.6 for convex domains with holes of identical size, for simplicity. Analogous lower bounds in the generic setting of convex domains with multiple holes of varying size are presented in Theorem 5.3. For $p\in\{2,\dots,n-1\}$, we further highlight the necessity of the partition parameter $R_{\scriptscriptstyle\hskip-0.73616pt\mathcal{P}}$ (see Section 5) in the lower bounds by constructing a family of convex domains with multiple holes (with the diameter of the domain and the radius of the largest hole uniformly bounded), satisfying the hypothesis of Theorem 5.3, having arbitrarily small eigenvalues as $R_{\scriptscriptstyle\hskip-0.73616pt\mathcal{P}}\to 0$. See Theorem 5.5.
The geometric assumption in Theorem 1.6 (and Theorem 5.3) may be removed, with the price that we obtain lower bounds for higher order exact eigenvalues instead of the smallest one. These lower bounds are later presented in Theorem 5.1 for convex domains with possibly non-identical holes. The exponent of the ratio $\hat{R}_{c}/D$ in these bounds is independent of the number of holes, in contrast to the expressions of Theorems 1.6, and instead, a factor of ${\mathfrak{h}}^{2p-1}$ appears in the denominator of the lower bounds.
The proof of Theorem 1.6 (resp. Theorem 5.3) is based on an induction approach where we consider the sequence of annular domains with convex outer boundary obtained from the Voronoi partition (resp. Laguerre-Voronoi partition) of $\mathfrak{D}$ given by the hypothesis and use an adaptation of the gluing lemmas of McGowan [McG93, Lemma 2.3] and Gentile-Pagliara [GP95, Lemma 1] to recursively employ the lower bounds of Theorem 1.1 for the eigenvalues of the pieces and the earlier mentioned bounds of [Gue04b] for the convex intersections of their open neighborhoods constructed suitably. On the other hand, we prove Theorem 5.1 by generalising McGowan’s lemma to higher degree forms under appropriate topological assumptions on the open cover and applying it to the cover of the domain $\mathfrak{D}$ obtained from its Laguerre-Voronoi partition, in combination with the bounds mentioned above. See Section 5 for the details.
This article is organised as follows. In Section 2 on preliminaries, we introduce the absolute and the relative eigenvalue problems for the Hodge Laplacian on compact Riemannian manifolds with boundary, followed by a brief description of the standard spectral properties and a version of the Hodge decomposition that is of relevance to us. In Section 3, we prove the lower bounds of Theorem 1.1 for the first exact eigenvalue of annular domains, and Theorem 1.3 for the lower bounds on the Neumann eigenvalues of a manifold in terms of those that of the covering subdomains. In Section 4, we prove Theorem 1.5 by constructing annular domains with arbitrarily small eigenvalues and emphasise the necessity of the contact radius in our lower bounds. In Section 5, we describe the Laguerre-Voronoi partition for domains with multiple holes, then state and prove Theorem 5.1 about the lower bounds for higher order exact eigenvalues of domains with multiple holes, and Theorem 5.3 that generalises Theorem 1.6 dealing with the lower bounds for the smallest exact eigenvalue for convex domains with multiple holes. We also recall a few versions of McGowan’s lemma as part of the proofs and present a generalisation of it to higher degree forms, the proof of which is somewhat technical and is the subject of the Appendix.
2. Preliminaries
Let $(M^{n},g)$ be a compact, connected, orientable, smooth Riemannian $n$-manifold with boundary. As mentioned in the introduction, the Hodge Laplacian on the space of differential $p$-forms $\Omega^{p}(M)$, for $p\in\{0,\dots,n\}$, is defined by
$\Delta^{(p)}$ $:\Omega^{p}(M)\to\Omega^{p}(M)$
$\omega$ $\mapsto(d\delta+\delta d)\omega,$
where $d$ and $\delta$ denote the exterior derivative and the codifferential, respectively, with the convention that $d\equiv 0$ on $n$-forms and $\delta\equiv 0$ on functions. We use the standard notation $\langle.\,.\rangle$ to denote the canonical pointwise inner product of $p$-forms induced by the Riemannian metric, and $\langle\hskip-3.0pt\langle{.},{.}\rangle\hskip-3.0pt\rangle_{L^{2}\Omega^{p}(M)}$ to denote the $L^{2}$-innerproduct of square integrable forms on $M$; $|\alpha|$ and $\|\alpha\|$ to denote the norms given by $\langle\alpha,\alpha\rangle^{\frac{1}{2}}$ and $\langle\hskip-3.0pt\langle{\alpha},{\alpha}\rangle\hskip-3.0pt\rangle_{L^{2}\Omega^{p}(M)}^{\frac{1}{2}}$, respectively. The notation $\iota^{*}$ and $i_{\mathpzc{n}}$ stand for the pullback of a differential form with respect to the inclusion map $\iota:\partial M\hookrightarrow M$, and the interior product with respect to the unit outward normal vector field $\mathpzc{n}$} on the boundary.
Absolute eigenvalue problem: For $\omega\in\Omega^{p}(M)$,
$\begin{cases}\Delta^{(p)}\omega=\mu\,\omega&\text{in $M$},\\ i_{\mathpzc{n}}\omega=0\text{ and }i_{\mathpzc{n}}d\omega=0&\text{on $\partial M$},\end{cases}$
where $\mu$ is the spectral parameter. The Hodge Laplacian with absolute boundary conditions, which generalises the Neumann problem for functions, has a discrete spectrum
$0=\mu^{(p)}_{0,1}=\dots=\mu^{(p)}_{0,b_{p}}<\mu^{(p)}_{1}\leq\mu^{(p)}_{2}\leq\dots\nearrow$$\infty$
with finite multiplicities. Here, the dimension of the kernel, $b_{p}\in\{0\}\cup\mathbb{N}$, is the absolute Betti number of $M$ (defined to be the rank of the $p^{\text{th}}$ homology group of $M$), a purely topological quantity independent of the metric. The $k^{\text{th}}$ positive eigenvalue has the variational characterisation
$\mu^{(p)}_{k}(M)=\min_{V_{b_{p}+k}}\,\,\max_{\omega\in V_{b_{p}+k}\setminus\{0\}}\frac{\|d\omega\|_{L^{2}\Omega^{p+1}(M)}^{2}+\|\delta\omega\|_{L^{2}\Omega^{p-1}(M)}^{2}}{\|\omega\|_{L^{2}\Omega^{p}(M)}^{2}},$
where $V_{b_{p}+k}$ denotes a $(b_{p}+k)$-dimensional subspace of $p$-forms with finite $H^{1}$-norm that are tangential to the boundary (i.e., $i_{\mathpzc{n}}\omega=0$). We refer to the text [Sch95] for the definition of Sobolev norms on forms.
We recall a version of the Hodge decomposition for $p$-forms [Tay11, Proposition 9.8] that is of relevance to the absolute eigenvalue problem. Given $\omega\in\Omega^{p}(M)$, there exist an $\alpha\in\Omega^{p}(M)$ and a harmonic form $\gamma\in\Omega^{p}(M)$ satisfying the absolute boundary conditions, $i_{\mathpzc{n}}\alpha=0$, $i_{\mathpzc{n}}d\alpha=0$, $i_{\mathpzc{n}}\gamma=0$ and $i_{\mathpzc{n}}d\gamma=0$, such that
$\omega=d\delta\alpha+\delta d\alpha+\gamma,$ (2.1)
and the three terms on the RHS are orthogonal in $L^{2}\Omega^{p}(M)$. Note that $\gamma$, being a harmonic form satisfying the absolute boundary conditions, is closed and coclosed. As is well-known, it follows from this decomposition that the spectrum of the Hodge Laplacian splits into two families, the exact $p$-spectrum
$0<\mu^{{(p)}}_{{1,\operatorname{\mathrm{ex}}}}\leq\mu^{{(p)}}_{{2,\operatorname{\mathrm{ex}}}}\leq\dots\nearrow$$\infty$,
and the coexact $p$-spectrum
$0<\mu^{(p)}_{1,\operatorname{\mathrm{co}}}\leq\mu^{(p)}_{2,\operatorname{\mathrm{co}}}\leq\dots\nearrow$$\infty$,
in addition to the eigenvalue zero with finite (possibly zero) multiplicity. Moreover, the coexact $p$-spectrum is identical to the exact $(p+1)$-spectrum, for $p\in\{0,\dots,n-1\}$:
$\mu^{(p)}_{k,\operatorname{\mathrm{co}}}=\mu^{{(p+1)}}_{{k,\operatorname{\mathrm{ex}}}},\qquad\forall\,k\in\mathbb{N}.$
These spectra satisfy the variational characterisations given by
$\mu^{{(p)}}_{{k,\operatorname{\mathrm{ex}}}}(M)=\min_{V_{k}^{\operatorname{\mathrm{ex}}}}\max_{\omega\in V_{k}^{\operatorname{\mathrm{ex}}}\setminus\{0\}}\frac{\|\delta\omega\|_{L^{2}\Omega^{p-1}(M)}^{2}}{\|\omega\|_{L^{2}\Omega^{p}(M)}^{2}}\quad\text{and}\quad\mu^{(p)}_{k,\operatorname{\mathrm{co}}}(M)=\min_{V_{k}^{\operatorname{\mathrm{co}}}}\max_{\omega\in V_{k}^{\operatorname{\mathrm{co}}}\setminus\{0\}}\frac{\|d\omega\|_{L^{2}\Omega^{p+1}(M)}^{2}}{\|\omega\|_{L^{2}\Omega^{p}(M)}^{2}},$
where $V_{k}^{\operatorname{\mathrm{ex}}}$ (resp. $V_{k}^{\operatorname{\mathrm{co}}}$) is a $k$-dimensional subspace of the exact (resp. coexact) $p$-forms with finite $H^{1}$-norm that are tangential to the boundary. The above characterisations also hold with $V_{k}^{\operatorname{\mathrm{ex}}}$ (resp. $V_{k}^{\operatorname{\mathrm{co}}}$) replaced by $V_{b_{p}+k}^{\operatorname{cl}}$ (resp. $V_{b_{p}+k}^{\operatorname{cc}}$), a $(b_{p}+k)$-dimensional subspace of the closed (resp. coclosed) $p$-forms with finite $H^{1}$-norm that are tangential to the boundary.
Furthermore, the following variant of the variational characterisation for the exact $p$-spectrum (see, for example, [Dod82, Proposition 3.1]) would be of use in our proofs, as it has no restrictions on the boundary conditions for the test forms. For $p\in\{1,\dots,n\}$,
$\mu^{{(p)}}_{{k,\operatorname{\mathrm{ex}}}}(M)=\inf_{V_{k}}\sup_{\eta\in V_{k}\setminus\{0\}}\,\sup_{\theta\,:\,\eta=d\theta}\frac{\|\eta\|^{2}_{L^{2}\Omega^{p}(M)}}{\|\theta\|^{2}_{L^{2}\Omega^{p-1}(M)}},$ (2.2)
where $V_{k}$ ranges over all $k$-dimensional subspaces of the space of $L^{2}$-integrable exact $p$-forms on $M$.
It is a natural choice to use the absolute boundary conditions in the study of the Hodge Laplacian spectrum for manifolds with boundary, but the following dual version of this problem, known as the relative eigenvalue problem that generalises the Dirichlet problem for functions, is also of significance: For $\omega\in\Omega^{p}(M)$,
$\begin{cases}\Delta^{(p)}\omega=\lambda\omega&\text{in $M$},\\ \iota^{*}\omega=0\text{ and }\iota^{*}\delta\omega=0&\text{on $\partial M$},\end{cases}$
where $\lambda$ is the spectral parameter. It has a discrete spectrum
$0=\lambda^{(p)}_{0,1}=\dots=\lambda^{(p)}_{0,b_{p}^{\operatorname{rel}}}<\lambda^{(p)}_{1}\leq\lambda^{(p)}_{2}\leq\dots\nearrow$$\infty$
with finite multiplicities. The dimension of the kernel, $b_{p}^{\operatorname{rel}}\in\{0\}\cup\mathbb{N}$, is the relative Betti number of $(M,\partial M)$. The absolute and the relative eigenvalue problems are dual to each other and are connected by the Hodge star operator. Their spectra obey, for $p\in\{0,\dots,n\}$,
$\mu^{(p)}_{b_{p}+k}=\lambda^{(n-p)}_{b_{p}^{\operatorname{rel}}+k},\qquad\forall\,k\in\mathbb{N}.$
The relative eigenvalue problem has analogous variational characterisation, and a spectral decomposition into the exact, coexact and harmonic forms with appropriate boundary conditions that can be deduced from a dual version of the Hodge decomposition stated above. Naturally, the lower bounds stated in this article for the absolute eigenvalues have counterparts for the relative eigenvalues.
3. Lower bounds for the smallest positive eigenvalue of convex domains with a hole
In this section, we prove Theorem 1.1 and Theorem 1.3. Let $\mathfrak{D}={\widetilde{\mathfrak{D}}}\setminus\overline{\mathbb{B}_{R_{h}}}$ be an annular domain in $\mathbb{R}^{n}$, where $\widetilde{\mathfrak{D}}$} is a convex domain in $\mathbb{R}^{n}$ and $\mathbb{B}_{R_{h}}\subsetneq{\widetilde{\mathfrak{D}}}$ is a Euclidean $n$-ball of radius $R_{h}$. Let $R_{c}$ and $D$ be the contact radius and the diameter of $\mathfrak{D}$, and $\mathbb{S}_{R_{h}}=\partial\mathbb{B}_{R_{h}}$ denote the round sphere of radius $R_{h}$. Fix $n\geq 2$ and $p\in\{1,\dots,n-1\}$. We divide the proof into two cases– when $R_{h}\geq\frac{R_{c}}{2}$, and when $R_{h}<\frac{R_{c}}{2}$.
Case 1: $R_{h}\geq\frac{R_{c}}{2}$.
[…]
Lemma 3.1.
For a compact Riemannian manifold $(M,g)$ with boundary, if $\tilde{\omega}\in\Omega^{p}(M)$ is closed and coexact (say, $\tilde{\omega}=\delta\hat{\omega}$) with $i_{\mathpzc{n}}\hat{\omega}=0$, then $\tilde{\omega}=0$.
Proof.
From the Hodge decomposition (2.1), we may write
$\tilde{\omega}=d\tilde{\beta}+\delta\tilde{\alpha}+\tilde{h}$,
where $\tilde{\alpha}$ is an exact form with $i_{\mathpzc{n}}\alpha=0$ and $\tilde{h}$ is a harmonic form satisfying the absolute boundary conditions (hence closed and coclosed). As $\tilde{\omega}$ is closed, we have that $d\delta\tilde{\alpha}=0$, which, using Green’s identity, implies
$\langle\hskip-3.0pt\langle{\delta\tilde{\alpha}},{\delta\tilde{\alpha}}\rangle\hskip-3.0pt\rangle_{L^{2}\Omega^{p}(M)}=\langle\hskip-3.0pt\langle{d\delta\tilde{\alpha}},{\tilde{\alpha}}\rangle\hskip-3.0pt\rangle_{L^{2}\Omega^{p}(M)}-\langle\hskip-3.0pt\langle{\iota^{*}\delta\tilde{\alpha}},{i_{\mathpzc{n}}d\tilde{\alpha}}\rangle\hskip-3.0pt\rangle_{L^{2}\Omega^{p}(\partial M)}=0,$
hence the Hodge decomposition of $\tilde{\omega}$ has no $\delta\tilde{\alpha}$ component. Now,
$\langle\hskip-3.0pt\langle{\omega},{\omega}\rangle\hskip-3.0pt\rangle_{L^{2}\Omega^{p}(M)}$ $=\langle\hskip-3.0pt\langle{\delta\hat{\omega}},{d\tilde{\beta}+\tilde{h}}\rangle\hskip-3.0pt\rangle_{L^{2}\Omega^{p}(M)}$
$=\langle\hskip-3.0pt\langle{\delta\hat{\omega}},{d\tilde{\beta}}\rangle\hskip-3.0pt\rangle_{L^{2}\Omega^{p}(M)}+\langle\hskip-3.0pt\langle{\delta\hat{\omega}},{\tilde{h}}\rangle\hskip-3.0pt\rangle_{L^{2}\Omega^{p}(M)}$
$=\langle\hskip-3.0pt\langle{\delta\delta\hat{\omega}},{\tilde{\beta}}\rangle\hskip-3.0pt\rangle_{L^{2}\Omega^{p}(M)}+\langle\hskip-3.0pt\langle{\iota^{*}\tilde{\beta}},{i_{\mathpzc{n}}\delta\hat{\omega}}\rangle\hskip-3.0pt\rangle_{L^{2}\Omega^{p}(\partial M)}+\langle\hskip-3.0pt\langle{\hat{\omega}},{d\tilde{h}}\rangle\hskip-3.0pt\rangle_{L^{2}\Omega^{p}(M)}-\langle\hskip-3.0pt\langle{\iota^{*}\tilde{h}},{i_{\mathpzc{n}}\hat{\omega}}\rangle\hskip-3.0pt\rangle_{L^{2}\Omega^{p}(\partial M)}$
$=0.$
Hence,
$\mu^{{(p)}}_{{1,\operatorname{\mathrm{ex}}}}(\mathfrak{D})=\frac{\|d\alpha\|^{2}_{L^{2}\Omega^{p}(\mathfrak{D})}}{\|\alpha\|^{2}_{L^{2}\Omega^{p-1}(\mathfrak{D})}}.$
It is difficult to directly find explicit geometric lower bounds for the term on the RHS, but we observe that $\alpha$ minimises the $L^{2}$-norm among all $\eta\in\Omega^{p-1}(\mathfrak{D})$ such that $d\eta=\omega$: If $\eta$ is any such form, then $\eta-\alpha$ is closed, so from the Hodge decomposition (2.1), there exists a harmonic form $h$ satisfying the absolute boundary conditions $i_{\mathpzc{n}}h=0$ and $i_{\mathpzc{n}}dh=0$, and an exact form $\beta$ with appropriate boundary conditions, such that $\eta-\alpha=\beta+h$. Further, $\alpha$, $\beta$ and $h$ would be mutually orthogonal, and we have
$\|\alpha\|_{L^{2}\Omega^{p-1}(\mathfrak{D}))}\leq\|\eta\|_{L^{2}\Omega^{p-1}(\mathfrak{D}))}.$
This implies that
$\mu^{{(p)}}_{{1,\operatorname{\mathrm{ex}}}}(\mathfrak{D})\geq\frac{\|\omega\|^{2}_{L^{2}\Omega^{p}(\mathfrak{D}))}}{\|\eta\|^{2}_{L^{2}\Omega^{p-1}(\mathfrak{D}))}},$ (3.1)
giving us the flexibility in terms of the choice of $\eta$. We adapt the methods from [CT97] to construct a primitive $\eta\in\Omega^{p-1}(\mathfrak{D})$ (i.e., $d\eta=\omega$) such that its $L^{2}$-norm is controlled by that of $\omega$ explicitly in terms of the geometry of $\mathfrak{D}$. More precisely, we prove the following:
Proposition 3.2.
There exists a primitive $\eta\in\Omega^{p-1}(\mathfrak{D})$ of $\omega$ such that
$\frac{\|\eta\|_{L^{2}\Omega^{p-1}(\mathfrak{D})}}{\|\omega\|_{L^{2}\Omega^{p}(\mathfrak{D})}}\leq KR_{c}\left(\frac{D}{R_{c}}\right)^{(n+1)p+\frac{n}{2}}\left({\frac{R_{h}}{R_{c}}}\right)^{(n-1)(\frac{3p}{2}+2)+p}.$
We would then conclude from (3.1) that, when $R_{h}\geq\frac{R_{c}}{2}$
$\mu^{{(p)}}_{{1,\operatorname{\mathrm{ex}}}}(\mathfrak{D})\geq\frac{\|\omega\|^{2}_{L^{2}\Omega^{p}(\mathfrak{D})}}{\|\eta\|^{2}_{L^{2}\Omega^{p-1}(\mathfrak{D})}}\geq K\frac{1}{R_{c}^{2}}\left(\frac{R_{c}}{D}\right)^{2(n+1)p+n}\left(\frac{R_{c}}{R_{h}}\right)^{(3n-1)p+4(n-1)}.$
Let us proceed with the proof of Proposition 3.2. We utilise the following estimate for convex domains, by applying it to sufficiently small pieces of the annular domain that are convexified with a control on the volumes and the diameters. We then use these local estimates to deduce a global Poincaré-type inequality, using an explicit isomorphism between Cěch and de Rham cohomology.
Lemma 3.3.
[CT97, Lemma 2.1] For a bounded, convex domain $\mathcal{D}\in\mathbb{R}^{n}$, there exists a constant $K(n)>0$ such that, given any closed form $\tilde{\omega}\in\Omega^{p}(\overline{\mathcal{D}})$, for $p\in\{1,\dots,n\}$, there exists a solution $\tilde{\phi}\in\Omega^{p-1}(\overline{\mathcal{D}})$ to the equation $d\tilde{\phi}=\tilde{\omega}$ that satisfies
$\|\tilde{\phi}\|_{L^{2}\Omega^{p-1}(\mathcal{D})}\leq K\frac{\operatorname{\mathrm{Diam}}(\mathcal{D})^{n+1}}{\operatorname{\mathrm{Vol}}(\mathcal{D})}\|\tilde{\omega}\|_{L^{2}\Omega^{p}(\mathcal{D})}.$
The proof of this lemma from [CT97] also holds for domains which are piecewise smooth.
Figure 3.
For convenience, we choose the centre of $\mathbb{S}_{R_{h}}$ to be the origin. Define
$r_{0}\coloneqq KR_{c},$
for sufficiently small $K$ described later by (3.6). Choose an open cover of $\mathbb{S}_{R_{h}}$ given by balls $\{\mathcal{B}_{i}\}_{i=1}^{N}$ of radii $\frac{r_{0}}{4^{n}}$ on $\mathbb{S}_{R_{h}}$, with their centers $\{t_{i}\}_{i=1}^{N}$ at least $\frac{r_{0}}{4^{n}}$ apart mutually with respect to the round metric. Observe that
$N\leq\frac{\operatorname{\mathrm{Vol}}(\mathbb{S}_{R_{h}})}{\operatorname{\mathrm{Vol}}(\frac{1}{2}\mathcal{B}_{i})}\leq K\left(\frac{R_{h}}{r_{0}}\right)^{n-1}\leq K\left({\frac{R_{h}}{R_{c}}}\right)^{n-1},$ (3.2)
where $\frac{1}{2}\mathcal{B}_{i}$ denotes the ball on $\mathbb{S}_{R_{h}}$ concentric to $\mathcal{B}_{i}$ with half its radius. Now, $\{U_{i}\coloneqq\mathfrak{D}\cap\mathpzc{P}^{-1}(\mathcal{B}_{i})\}_{i=1}^{N}$ will be an open cover for $\mathfrak{D}$, where $\mathpzc{P}$ is the orthogonal projection from $\mathbb{R}^{n}$ onto $\mathbb{S}_{R_{h}}$. Our choice of $r_{0}$ above also ensures that $\{\mathcal{B}_{i}\}$ are geodesically convex, and so are all the nontrivial intersections $\cap_{k=1}^{k_{0}}\mathcal{B}_{i_{k}}$, for $k_{0}\in\{1,\dots,N\}$. It then follows that $\{U_{i}\}_{i=1}^{n}$ is a good cover of $\mathfrak{D}$, i.e., all the nontrivial intersections $\cap_{k=1}^{k_{0}}U_{i_{k}}$ are contractible for all $k_{0}\in\{1,\dots,N\}$. We use this property later while invoking the Čech cohomology. Furthermore, we have, for $i\in\{1,\dots,N\}$,
$KD^{n}\geq\operatorname{\mathrm{Vol}}(U_{i})\geq KR_{c}r_{0}^{n-1}\geq KR_{c}^{n}.$ (3.3)
For $q\in\{0,\dots,N-1\}$ and a multi-index $I=(i_{0},\dots,i_{q})$ such that $1\leq i_{0}<\dots<i_{q}\leq N$, denote by $U_{I}$ the intersection $\cap_{k=0}^{q}U_{i_{k}}$. Let $S_{q}$ be the set of such multi-indices of length $q+1$ with $U_{I}\neq\varnothing$. Note that
$\operatorname{\mathrm{Card}}(S_{q})\leq N(N-1)\dots(N-q)\leq N^{q+1}.$
Define, for $\lambda\in\{0,\dots,n\}$,
$U_{i}^{\lambda}\coloneqq\mathfrak{D}\cap\mathpzc{P}^{-1}\left(\mathcal{B}\Big(t_{i},\frac{4^{\lambda}}{4^{n}}r_{0}\Big)\right),$
and
$U_{I}^{\lambda}\coloneqq U_{i_{0}}^{\lambda}\cap\dots\cap U_{i_{q}}^{\lambda},$
for each $I=(i_{0},\dots,i_{q})\in S_{q}$. We have that, for $\lambda\in\{0,\dots,n-1\}$,
$U_{i_{k}}^{\lambda}\subset U_{I}^{\lambda+1},\qquad\forall\,k\in\{0,\dots,q\},\forall\,I\in S_{q}.$
Let $\{\tilde{\varphi}_{i}\}_{i=1}^{N}$ be a partition of unity subordinate to the cover $\{\mathcal{B}_{i}\}_{i=1}^{N}$ of $\mathbb{S}_{R_{h}}$ such that $\|d\tilde{\varphi}_{i}\|_{\infty}\leq\frac{K}{r_{0}}$ for all $i\in\{1,\dots,N\}$, for some constant $K(n)$. This gives rise to a partition of unity $\{\varphi_{i}\coloneqq\tilde{\varphi}_{i}\circ\mathpzc{P}\}_{i=1}^{N}$ subordinate to the cover $\{U_{i}\}_{i=1}^{N}$ of $\mathfrak{D}$, also satisfying the bound $\|d\varphi_{i}\|_{\infty}\leq\frac{K}{R_{c}}$ for all $i\in\{1,\dots,N\}$.
Lemma 3.4.
Let $\lambda\in\{1,\dots,n\}$. There exists a constant $K(n)$ such that, given an exact form $\tilde{\omega}\in\Omega^{p}\left(\overline{U_{I}^{\lambda}}\right)$, there exists $\tilde{\phi}\in\Omega^{p}\left(\overline{U_{I}^{\lambda}}\right)$ such that $d\tilde{\phi}=\tilde{\omega}$ on $U_{I}^{\lambda}$ and
$\|\tilde{\phi}\|_{L^{2}\Omega^{p-1}(U_{I}^{\lambda})}\leq K\frac{D^{n+1}}{R_{c}^{n}}\|\tilde{\omega}\|_{L^{2}\Omega^{p}(U_{I}^{\lambda})}.$ (3.4)
Proof.
For $t\in\mathbb{S}_{R_{h}}$, set
[…]
Figure 4.
We construct a quasi-isometry as follows, that transforms a sufficiently small piece of $\mathfrak{D}$ into a convex domain with control on the Jacobian of this transformation. Choose a coordinate system whose origin is at the centre of $\mathbb{B}_{R_{h}}$ and such that the point $t$ corresponds to $(R_{h},0,\dots,0)$. Define
$\Phi:\mathcal{O}_{\theta}$ $\to\mathbb{R}^{n}$
$(x_{1},\dots,x_{n})$ $\mapsto\frac{R\cos\theta}{x_{1}}(x_{1},\dots,x_{n}),$
where $R=\sqrt{\sum_{i=1}^{n}x_{i}^{2}}$. Let $\alpha:[R_{h},\infty)\to[0,1]$ be a smooth cutoff function that is equal to $1$ on $R_{h},R_{h}+\frac{R_{c}}{3}]$ and vanishing on $R_{h}+\frac{2R_{c}}{3},\infty)$, whose gradient is controlled by $|\alpha^{\prime}|\leq\frac{K}{R_{c}}$. The required quasi-isometry is then given by
$\Psi(x_{1},\dots,x_{n})\coloneqq\alpha(R)\Phi(x_{1},\dots,x_{n})+(1-\alpha(R))(x_{1},\dots,x_{n}).$
We have, for $k,l\in\{1,\dots,n\}$,
$\frac{\partial\Psi_{k}}{\partial x_{l}}=\alpha(R)\frac{\partial\Phi_{k}}{\partial x_{l}}(x_{1},\dots,x_{n})+\alpha^{\prime}(R)\frac{x_{l}}{R}\Phi_{k}(x_{1},\dots,x_{n})+(1-\alpha(R))\delta_{kl}-\alpha^{\prime}(R)\frac{x_{l}x_{k}}{R}.$
Substituting the expression for $\Phi$ and using that
$x_{1}$ $=R\cos(\theta(x))\geq R\cos\theta,$
$|x_{i}|$ $\leq R\sin(\theta(x))\leq R\sin\theta,\qquad\forall\,i\in\{2,\dots,n\},$
$\alpha$ $=\alpha^{\prime}\equiv 0\qquad\text{on }\left[R_{h}+\frac{2R_{c}}{3},\infty\right),$
one can deduce that
[…]
By the equivalence of the max norm and the operator norm for matrices, let $K_{1}(n)\geq 1$ be a constant such that
$K_{1}^{-1}\|.\|_{\max}\leq\|.\|_{\operatorname{op}}\leq K_{1}\|.\|_{\max}.$
Set
$\theta_{0}\coloneqq\frac{1}{3K_{1}K_{0}{\frac{R_{h}}{R_{c}}}},$
requiring without loss of generality that the RHS is less than $\frac{\pi}{2}$. Then, for $\theta\in(0,\theta_{0}]$,
[…]
and
[…]
$\leq K_{1}K_{0}{\frac{R_{h}}{R_{c}}}\theta+\left(K_{1}K_{0}{\frac{R_{h}}{R_{c}}}\theta\right)^{2}+\dots$
$=\frac{K_{1}K_{0}{\frac{R_{h}}{R_{c}}}\theta}{1-K_{1}K_{0}{\frac{R_{h}}{R_{c}}}\theta}$
$\overset{\text{using }\theta\leq\theta_{0}}{\leq}\frac{1}{2}.$
Thus,
[…]
Set
$r_{0}\coloneqq R_{h}\theta_{0}=KR_{c},$ (3.6)
so that $r_{0}<\frac{\pi}{2}R_{h}$ and the balls $\mathcal{B}(t_{i},r_{0})$ are geodesically convex on $\mathbb{S}_{R_{h}}$.
Let $(U_{I}^{\lambda})^{\prime}\subset\mathbb{R}^{n}$ be the image of $U_{I}^{\lambda}\subset\mathcal{O}_{\theta_{0}}$ under the diffeomorphism $\Psi$. We have
$\operatorname{\mathrm{Vol}}(U_{I}^{\lambda})$ $=\int_{U_{I}^{\lambda}}\operatorname{d\hskip-1.00006ptV}$
$=\int_{(U_{I}^{\lambda})^{\prime}}|\det(\operatorname{\mathrm{Jac}}\Psi^{-1})|\,\operatorname{d\hskip-1.00006ptV}$
$\leq\|\operatorname{\mathrm{Jac}}\Psi^{-1}\|_{\operatorname{op}}^{n}\int_{(U_{I}^{\lambda})^{\prime}}\operatorname{d\hskip-1.00006ptV}$
[…]
We also have that $\operatorname{\mathrm{Diam}}((U_{I}^{\lambda})^{\prime})=\operatorname{\mathrm{Diam}}(U_{I}^{\lambda})$, since $(U_{I}^{\lambda})^{\prime}$ is in fact the convex-hull of $U_{I}^{\lambda}$.
We recall [CT97, Lemma 2.1]: For a bounded, convex domain $\mathcal{D}\in\mathbb{R}^{n}$, there exists a constant $K(n)>0$ such that, given any closed form $\tilde{\omega}\in\Omega^{p}(\overline{\mathcal{D}})$, for $p\in\{1,\dots,n\}$, there exists a solution $\tilde{\phi}\in\Omega^{p-1}(\overline{\mathcal{D}})$ to the equation $d\tilde{\phi}=\tilde{\omega}$ that satisfies
$\|\tilde{\phi}\|_{L^{2}\Omega^{p-1}(\mathcal{D})}\leq K\frac{\operatorname{\mathrm{Diam}}(\mathcal{D})^{n+1}}{\operatorname{\mathrm{Vol}}(\mathcal{D})}\|\tilde{\omega}\|_{L^{2}\Omega^{p}(\mathcal{D})}.$
Applying this to the exact form $(\Psi^{-1})^{*}\tilde{\omega}$, the pullback of $\tilde{\omega}$ with respect to $\Psi^{-1}:(U_{I}^{\lambda})^{\prime}\to U_{I}^{\lambda}$, there exists $\hat{\phi}\in\Omega^{p-1}((U_{I}^{\lambda})^{\prime})$ with $d\hat{\phi}=(\Psi^{-1})^{*}\tilde{\omega}$ such that
$\left(\int_{(U_{I}^{\lambda})^{\prime}}|\hat{\phi}|^{2}\,\operatorname{d\hskip-1.00006ptV}\right)^{\frac{1}{2}}\leq K\frac{\operatorname{\mathrm{Diam}}((U_{I}^{\lambda})^{\prime})^{n+1}}{\operatorname{\mathrm{Vol}}((U_{I}^{\lambda})^{\prime})}\left(\int_{(U_{I}^{\lambda})^{\prime}}|(\Psi^{-1})^{*}\tilde{\omega}|^{2}\,\operatorname{d\hskip-1.00006ptV}\right)^{\frac{1}{2}}.$
Choose $\tilde{\phi}=\Psi^{*}\hat{\phi}$. Using diameter and volume comparison from before, in addition to the control on the Jacobian, we get
$\left(\int_{U_{I}^{\lambda}}|\tilde{\phi}|^{2}\,\operatorname{d\hskip-1.00006ptV}\right)^{\frac{1}{2}}$ $=\left(\int_{(U_{I}^{\lambda})^{\prime}}|\det\operatorname{\mathrm{Jac}}\Psi^{-1}|\,|(\Psi^{-1})^{*}\tilde{\phi}|^{2}\,\operatorname{d\hskip-1.00006ptV}\right)^{\frac{1}{2}}$
$\leq\left(\frac{3}{2}\right)^{\frac{n}{2}}\left(\int_{(U_{I}^{\lambda})^{\prime}}|\hat{\phi}|^{2}\,\operatorname{d\hskip-1.00006ptV}\right)^{\frac{1}{2}}$
$\leq K\frac{\operatorname{\mathrm{Diam}}((U_{I}^{\lambda})^{\prime})^{n+1}}{\operatorname{\mathrm{Vol}}((U_{I}^{\lambda})^{\prime})}\left(\int_{(U_{I}^{\lambda})^{\prime}}|(\Psi^{-1})^{*}\tilde{\omega}|^{2}\,\operatorname{d\hskip-1.00006ptV}\right)^{\frac{1}{2}}$
$\leq K\frac{\operatorname{\mathrm{Diam}}(U_{I}^{\lambda})^{n+1}}{\operatorname{\mathrm{Vol}}(U_{I}^{\lambda})}\left(\int_{U_{I}^{\lambda}}|\det\operatorname{\mathrm{Jac}}\Psi|\,|\tilde{\omega}|^{2}\,\operatorname{d\hskip-1.00006ptV}\right)^{\frac{1}{2}}$
$\leq K\frac{\operatorname{\mathrm{Diam}}(\mathfrak{D})^{n+1}}{R_{c}^{n}}\left(\int_{U_{I}^{\lambda}}|\tilde{\omega}|^{2}\,\operatorname{d\hskip-1.00006ptV}\right)^{\frac{1}{2}},$
where we use that, for $\lambda\in\{1,\dots,n\}$,
[…]
Hence the lemma. ∎
Let ${\varepsilon}(i,I\setminus i)$ denote the sign of the permutation that orders $(i,I\setminus i)$.
Lemma 3.5.
There exists $\nu\in\Omega^{p-1}(\mathfrak{D})$, and, for every $I\in S_{p}$, a constant function $\eta_{I}$ on $U_{I}^{1}$ such that
$\omega-d\nu=-\sum_{I\in S_{p}}\left(\sum_{i\in I}{\varepsilon}(i,I\setminus i)\varphi_{i}\,d\varphi_{I\setminus i}\right)\eta_{I},$ (3.7)
with
$\|\nu\|_{L^{2}\Omega^{p-1}(\mathfrak{D})}\leq KR_{c}\left(\frac{D}{R_{c}}\right)^{(n+1)p}\left({\frac{R_{h}}{R_{c}}}\right)^{(n-1)p}\|\omega\|_{L^{2}\Omega^{p}(\mathfrak{D})},$ (3.8)
$\left(\operatorname{\mathrm{Vol}}(U_{I}^{1})\right)^{\frac{1}{2}}|\eta_{I}|=\|\eta_{I}\|_{L^{2}(\mathfrak{D})}\leq K\left(\frac{D^{n+1}}{R_{c}^{n}}\right)^{p}\|\omega\|_{L^{2}\Omega^{p}(\mathfrak{D})},\qquad\text{and}$ (3.9)
$\sum_{j\in J}{\varepsilon}(j,J\setminus j)\,\eta_{J\setminus j}=0,\qquad\forall\,J\in S_{p+1}.$ (3.10)
Here, $d\varphi_{K}$ denotes the $(k+1)$-form $d\varphi_{i_{0}}\wedge\dots\wedge d\varphi_{i_{k}}$ for a multi-index $K=(i_{0},\dots,i_{k})$.
Proof.
Adapting the construction from [CT97, Section 2], we prove the following property by induction: For any $q\in\{1,\dots,p\}$, there exists a closed form $\eta_{I}^{(q)}\in\Omega^{p-q}(U_{I}^{p-q+1})$ for each $I\in S_{q}$, and $\nu^{(q)}\in\Omega^{p-1}(\mathfrak{D})$ such that
$\omega-d\nu^{(q)}=-\sum_{I\in S_{q}}\left(\sum_{i\in I}{\varepsilon}(i,I\setminus i)\varphi_{i}d\varphi_{I\setminus i}\right)\wedge\eta_{I}^{(q)},$ (3.11)
$\|\nu^{(q)}\|_{L^{2}\Omega^{p-1}(\mathfrak{D})}\leq KR_{c}\left(\frac{D}{R_{c}}\right)^{(n+1)q}\left({\frac{R_{h}}{R_{c}}}\right)^{(n-1)q}\|\omega\|_{L^{2}\Omega^{p}(\mathfrak{D})},$ (3.12)
$\|\eta_{I}^{(q)}\|_{L^{2}\Omega^{p-q}(U^{p-q+1})}\leq K\left(\frac{D^{n+1}}{R_{c}^{n}}\right)^{q}\|\omega\|_{L^{2}\Omega^{p}(\mathfrak{D})},$ (3.13)
and
$\sum_{j\in J}{\varepsilon}(j,J\setminus j)\,\eta^{(q)}_{J\setminus j}=0,\qquad\forall\,J\in S_{q+1}.$ (3.14)
For the base case $q=1$, we proceed as follows. We have from (3.4) that there exists, for each $i\in\{1,\dots,N\}$, an $\eta^{(1)}_{i}\in\Omega^{p-1}(U_{i}^{p})$ such that $d\eta^{(1)}_{i}=\omega$ and
$\|\eta^{(1)}_{i}\|_{L^{2}\Omega^{p-1}(U_{i}^{p})}\leq K\frac{D^{n+1}}{R_{c}^{n}}\|\omega\|_{L^{2}\Omega^{p}(\mathfrak{D})}.$
We then set $\nu^{(1)}\coloneqq\sum_{i=1}^{N}\varphi_{i}\eta^{(1)}_{i}$ so that
$\|\nu^{(1)}\|_{L^{2}\Omega^{p-1}(\mathfrak{D})}$ $\leq KN\frac{D^{n+1}}{R_{c}^{n}}\|\omega\|_{L^{2}\Omega^{p}(\mathfrak{D})}$
[…]
Using that $\sum_{i=1}^{N}d\varphi_{i}=0$ on $\mathfrak{D}$, a simple computation from [CT97, Section 2] gives
$\omega-d\nu^{(1)}=-\sum_{(i,j)\in S_{1}}(\varphi_{i}d\varphi_{j}-\varphi_{j}d\varphi_{i})\wedge\eta^{(2)}_{ij}$,
where $\eta^{(2)}_{ij}\in\Omega^{p-1}(U_{ij}^{p})$ is the closed form defined to be $\eta^{(1)}_{i}-\eta^{(1)}_{j}$ for $i<j$, satisfying
$\left(\int_{U^{p}_{ij}}\left|\eta^{(2)}_{ij}\right|^{2}\,\operatorname{d\hskip-1.00006ptV}\right)^{\frac{1}{2}}$ $\leq\left(\int_{U^{p}_{ij}}\left|\eta^{(1)}_{i}\right|^{2}+\left|\eta^{(1)}_{j}\right|^{2}+2\langle\eta^{(1)}_{i},\eta^{(1)}_{j}\rangle\,\operatorname{d\hskip-1.00006ptV}\right)^{\frac{1}{2}}$
$\leq\left(\int_{U^{p}_{ij}}\left|\eta^{(1)}_{i}\right|^{2}+\left|\eta^{(1)}_{j}\right|^{2}+2\max\left\{\left|\eta^{(1)}_{i}\right|^{2},\left|\eta^{(1)}_{j}\right|^{2}\right\}\,\operatorname{d\hskip-1.00006ptV}\right)^{\frac{1}{2}}$
$\leq K\frac{D^{n+1}}{R_{c}^{n}}\|\omega\|_{L^{2}\Omega^{p}(\mathfrak{D})}.$
Furthermore, for $(i,j,k)\in S_{2}$, we have
$u_{jk}-u_{ik}+u_{ij}=u_{j}-u_{k}-u_{i}+u_{k}+u_{i}-u_{j}=0.$
Suppose that the induction hypothesis holds for some $q\in\{1,\dots,p-1\}$. We now show that this would imply that the property holds for $q+1$. Again, we have from (3.4) that for each $I\in S_{q}$, there exists $\nu^{(q)}_{I}\in\Omega^{p-q-1}(U^{p-q+1}_{I})$ such that $d\nu^{(q)}_{I}=\eta^{(q)}_{I}$ on $U^{p-q+1}_{I}$ and
$\|\nu^{(q)}_{I}\|_{L^{2}\Omega^{p-q-1}(U^{p-q+1}_{I})}\leq K\frac{D^{n+1}}{R_{c}^{n}}\|\eta^{(q)}_{I}\|_{L^{2}\Omega^{p-q}(U^{p-q+1}_{I})}\leq K\left(\frac{D^{n+1}}{R_{c}^{n}}\right)^{q+1}\|\omega\|_{L^{2}\Omega^{p}(\mathfrak{D})},$
where we use the induction hypothesis in the second inequality. Now, define the (p1)(p-1)-form on $\mathfrak{D}$,
$\nu^{(q+1)}\coloneqq\nu^{(q)}+\sum_{I\in S_{q}}\sum_{i\in I}{\varepsilon}(i,I\setminus i)\varphi_{i}\,d\varphi_{I\setminus i}\wedge\nu^{(q)}_{I}$
and for each $J\in S_{q+1}$, define the (pq1)(p-q-1)-form on $U^{p-q}_{j}$,
$\eta^{(q+1)}_{J}=-(q+1)\sum_{j\in J}{\varepsilon}(j,J\setminus j)\nu^{(q)}_{J\setminus j}$.
The forms $\{\eta^{(q+1)}_{J}\}$ are closed by property (3.14) for $q$, and their $L^{2}$-norms are bounded by
$\|\eta^{(q+1)}_{J}\|_{L^{2}\Omega^{p-q-1}(U_{j}^{p-q})}$ $\leq(q+1)\sum_{j\in J}\|\nu^{(q)}_{J\setminus j}\|_{L^{2}\Omega^{p-q-1}(U_{j}^{p-q})}$
$\leq(q+1)(q+2)K\left(\frac{D^{n+1}}{R_{c}^{n}}\right)^{q+1}\|\omega\|_{L^{2}\Omega^{p}(\mathfrak{D})}$
$\leq K\left(\frac{D^{n+1}}{R_{c}^{n}}\right)^{q+1}\|\omega\|_{L^{2}\Omega^{p}(\mathfrak{D})}.$
One can then repeat the computations from [CT97, Section 2] to show that
$\omega-d\nu^{(q+1)}=-\sum_{J\in S_{q+1}}\left(\sum_{j\in J}{\varepsilon}(j,J\setminus j)\varphi_{j}d\varphi_{J\setminus j}\right)\wedge\eta_{J}^{(q+1)}$
and
$\sum_{j\in J}{\varepsilon}(j,J\setminus j)\,\eta^{(q+1)}_{J\setminus j}=0,\qquad\forall\,J\in S_{q+2}.$
Lastly, since $\operatorname{Supp}(\varphi_{i}\,d\varphi_{I\setminus i})\subset U^{p-q+1}_{I}$, we get from the definition of $\nu^{(q+1)}$,
$\|\nu^{(q+1)}\|_{L^{2}\Omega^{p-1}(\mathfrak{D})}-\|\nu^{(q)}\|_{L^{2}\Omega^{p-1}(\mathfrak{D})}$ $\leq\sum_{I\in S_{q}}\sum_{i\in I}\|\varphi_{i}\,d\varphi_{I\setminus i}\|_{\infty}\|\nu^{(q)}_{I}\|_{L^{2}\Omega^{p-q-1}(U^{p-q+1}_{I})}$
$\leq K\operatorname{\mathrm{Card}}(S_{q})(q+1){\frac{1}{R_{c}^{q}}}\left(\frac{D^{n+1}}{R_{c}^{n}}\right)^{q+1}\|\omega\|_{L^{2}\Omega^{p}(\mathfrak{D})}$
$\leq KN^{q+1}{\frac{1}{R_{c}^{q}}}\left(\frac{D^{n+1}}{R_{c}^{n}}\right)^{q+1}\|\omega\|_{L^{2}\Omega^{p}(\mathfrak{D})}$
[…]
which implies that
$\|\nu^{(q+1)}\|_{L^{2}\Omega^{p-1}(\mathfrak{D})}$ $\leq\|\nu^{(q)}\|_{L^{2}\Omega^{p-1}(\mathfrak{D})}+KR_{c}\left(\frac{D}{R_{c}}\right)^{(n+1)(q+1)}\left({\frac{R_{h}}{R_{c}}}\right)^{(n-1)(q+1)}\|\omega\|_{L^{2}\Omega^{p}(\mathfrak{D})}$
$\leq KR_{c}\left(\frac{D}{R_{c}}\right)^{(n+1)q}\left({\frac{R_{h}}{R_{c}}}\right)^{(n-1)q}\left[1+\left(\frac{D}{R_{c}}\right)^{n+1}\left({\frac{R_{h}}{R_{c}}}\right)^{n-1}\right]\|\omega\|_{L^{2}\Omega^{p}(\mathfrak{D})}$
$\leq KR_{c}\left(\frac{D}{R_{c}}\right)^{(n+1)(q+1)}\left({\frac{R_{h}}{R_{c}}}\right)^{(n-1)(q+1)}\|\omega\|_{L^{2}\Omega^{p}(\mathfrak{D})},$
using that $\frac{D}{R_{c}}\geq 1$ and ${\frac{R_{h}}{R_{c}}}\geq\frac{1}{2}$. Thus, the lemma follows by induction, where we finally choose $\nu=\nu^{(p)}\in\Omega^{p-1}(\mathfrak{D})$ and the constant functions $\eta_{I}=\eta^{(p)}_{I}$ on $U_{I}^{1}$. ∎
Recall that
$\operatorname{\mathrm{Vol}}(U_{I}^{1})\geq\operatorname{\mathrm{Vol}}(U_{i})\geq KR_{c}^{n}.$
So, by (3.9),
$|\eta_{I}|\leq K\frac{D^{(n+1)p}}{R_{c}^{n(p+\frac{1}{2})}}\,\|\omega\|_{L^{2}\Omega^{p}(\mathfrak{D})}.$ (3.15)
Consider the family of constants $(\eta_{I})_{I\in S_{p}}$ as a Čech cochain with respect to the good cover $\{U_{i}\}_{i=1}^{N}$. In view of (3.10), it is a Cěch cocycle. It then follows from (3.7) that the Cěch cohomology class of this cocycle is same as the de Rham cohomology class of $\omega$ (see [BT82, Propostion 9.5]). As $\omega$ is exact, it further means that $(\eta_{I})_{I\in S_{p}}$ is in fact a coboundary. Using this and proceeding as in [CT97, Section 2], we get that
$d\eta=\omega,$
where $\eta$ is constructed as
$\eta=\nu+\sum_{H\in S_{p-1}}\sum_{i\in H}{\varepsilon}(i,I\setminus i)c_{H}\left(1+\frac{1}{p}-\frac{p+1}{2}\varphi_{i}-\sum_{j\notin H}\varphi_{j}\right)\varphi_{i}\,d\varphi_{H\setminus i}.$
Here, $c_{H}$ are constants which satisfy the bounds
$\sum_{H\in S_{p-1}}|c_{H}|$ $\leq KN^{\frac{p}{2}+1}\sum_{I\in S_{p}}|\eta_{I}|$
[…]
$\leq K\left({\frac{R_{h}}{R_{c}}}\right)^{(n-1)(\frac{3p}{2}+2)}\frac{D^{(n+1)p}}{R_{c}^{n(p+\frac{1}{2})}}\,\|\omega\|_{L^{2}\Omega^{p}(\mathfrak{D})}.$
Thus,
$\frac{\|\eta\|_{L^{2}\Omega^{p-1}(\mathfrak{D})}}{\|\omega\|_{L^{2}\Omega^{p}(\mathfrak{D})}}$ $\leq K\left(R_{c}\left(\frac{D}{R_{c}}\right)^{(n+1)p}\left({\frac{R_{h}}{R_{c}}}\right)^{(n-1)p}\right)$
$\hskip 100.00015pt+K\left(\left({\frac{R_{h}}{R_{c}}}\right)^{(n-1)(\frac{3p}{2}+2)}\frac{D^{(n+1)p}}{R_{c}^{n(p+\frac{1}{2})}}\left({\frac{R_{h}}{R_{c}}}\right)^{p}\frac{1}{R_{c}^{p-1}}D^{\frac{n}{2}}\right)$
$\leq KR_{c}\left(\frac{D}{R_{c}}\right)^{(n+1)p+\frac{n}{2}}\left({\frac{R_{h}}{R_{c}}}\right)^{(n-1)(\frac{3p}{2}+2)+p},$
which completes the proof of Proposition 3.2.
Case 2: $R_{h}<\frac{R_{c}}{2}$.
We need the following property which follows from a straightforward adaptation of a result of Anné and Colbois [AC93, Theorem 1.1] concerning the spectral convergence for perforated domains, to the framework of manifolds with boundary.
Proposition 3.6.
Let $M$ be a compact Riemannian $n$-manifold with boundary ($n\geq 2$) and $M_{\epsilon}$ be the manifold obtained by removing a finite number of non-overlapping geodesic balls of radius at most ${\epsilon}>0$, for $\epsilon$} smaller than the injectivity radius of $M$. Then,
$\lim_{{\epsilon}\to 0}\mu^{{(p)}}_{{1,\operatorname{\mathrm{ex}}}}(M_{\epsilon})=\mu^{{(p)}}_{{1,\operatorname{\mathrm{ex}}}}(M),\qquad\forall\,p\in\{0,\dots,n-1\}.$
If $\mathfrak{D}$ is an annulus whose outer radius is $R_{c}$ and inner radius at most $\frac{R_{c}}{2}$, consider the annulus $\mathfrak{D}^{\prime}$ of outer radius $1$ and inner radius at most $\frac{1}{2}$, obtained from scaling $\mathfrak{D}$ by a factor of $\frac{1}{R_{c}}$}. By Proposition 3.6 and continuity of eigenvalues with respect to the metric, we have
$\mu^{{(p)}}_{{1,\operatorname{\mathrm{ex}}}}(\mathfrak{D}^{\prime})\geq K,$
which implies that
$\mu^{{(p)}}_{{1,\operatorname{\mathrm{ex}}}}(\mathfrak{D})\geq\frac{K}{R_{c}^{2}}.$
When $\mathfrak{D}$ is a generic domain with $R_{h}<\frac{R_{c}}{2}$, we proceed as follows.
For $n\geq 3$ and $p\in\{2,\dots,n-1\}$, we use a version of [GP95, Lemma 1] which extends [McG93, Lemma 2.3] by Cheeger and McGowan (where it was stated for closed manifolds but the proof is essentially the same for absolute boundary conditions):
Lemma 3.7.
Let $(M^{n},g)$ be a compact Riemannian manifold with boundary $(n\geq 3)$ such that $M=U_{1}\cup U_{2}$ for some smooth subdomains $U_{1},U_{2}$ with $H^{p-1}(U_{1}\cap U_{2},\mathbb{R})=0$. Let $\{\rho_{1},\rho_{2}\}$ be a partition of unity subordinate to $\{U_{1},U_{2}\}$ with
$c_{\rho}\coloneqq\max\{\|\nabla\rho_{1}\|_{\infty}^{2},\|\nabla\rho_{2}\|_{\infty}^{2}\}.$
Then we have, for $p\in\{2,\dots,n-1\}$,
$\mu^{{(p)}}_{{1,\operatorname{\mathrm{ex}}}}(M)\geq\frac{K}{\left(\frac{1}{\mu^{{(p)}}_{{1,\operatorname{\mathrm{ex}}}}(U_{1})}+\frac{1}{\mu^{{(p)}}_{{1,\operatorname{\mathrm{ex}}}}(U_{2})}\right)\left(\frac{c_{\rho}}{\mu^{{(p-1)}}_{{1,\operatorname{\mathrm{ex}}}}(U_{1}\cap U_{2})}+1\right)}.$
We write $\mathfrak{D}=\mathfrak{D}_{1}\cup\mathfrak{D}_{2}$, where $\mathfrak{D}_{1}$ is the annular region whose outer boundary is the same as that of $\mathfrak{D}$ and the inner boundary is a sphere of radius $\frac{R_{c}}{2}$ (concentric to $\mathbb{S}_{R_{h}}$), and $\mathfrak{D}_{2}$ is the annular region of inner radius $R_{h}$ and outer radius $R_{c}$ (also concentric to $R_{c}$). Choose a partition of unity $\{\hat{\rho}_{1},\hat{\rho}_{2}\}$ subordinate to $\{\mathfrak{D}_{1},\mathfrak{D}_{2}\}$ such that
$c_{\rho}\coloneqq\max\{\|\nabla\hat{\rho}_{1}\|_{\infty}^{2},\|\nabla\hat{\rho}_{2}\|_{\infty}^{2}\}\leq\frac{K}{(R_{c}-R_{h})^{2}}\leq\frac{K}{(\frac{R_{c}}{2})^{2}}.$
Applying Lemma 3.7 and noting that $H^{p-1}(\mathfrak{D}_{1}\cap\mathfrak{D}_{2},\mathbb{R})=H^{p-1}({\mathbb{S}^{n-1}},\mathbb{R})=0$ for $n\geq 3$ and $p\in\{2,\dots,n-1\}$, we get,
$\mu^{{(p)}}_{{1,\operatorname{\mathrm{ex}}}}(\mathfrak{D})$ $\geq\frac{K}{\left(R_{c}^{2}\left(\frac{D}{R_{c}}\right)^{2(n+1)p+n}\left(\frac{R_{h}}{R_{c}}\right)^{(3n-1)p+4(n-1)}+R_{c}^{2}\right)\left(\frac{R_{c}^{-2}}{R_{c}^{-2}}+1\right)}$
$\geq K\frac{1}{R_{c}^{2}}\left(\frac{R_{c}}{D}\right)^{2(n+1)p+n}.$
For $n\geq 3$ and $p\in\{2,\dots,n-1\}$, we thus have, for any annular region,
$\mu^{{(p)}}_{{1,\operatorname{\mathrm{ex}}}}(\mathfrak{D})$ $\geq K\frac{1}{R_{c}^{2}}\left(\frac{R_{c}}{D}\right)^{2(n+1)p+n}\min\left\{1,\left(\frac{R_{c}}{R_{h}}\right)^{(3n-1)p+4(n-1)}\right\}$
$\geq K\frac{1}{R_{c}^{2}}\left(\frac{R_{c}}{D}\right)^{n(2n-1)}\min\left\{1,\left(\frac{R_{c}}{R_{h}}\right)^{3(n-1)(n+1)}\right\}.$
When $n\geq 2$ and $p=1$, for $R_{h}<\frac{R_{c}}{2}$, we take $\mathfrak{D}_{1}$ and $\mathfrak{D}_{2}$ as before and obtain from Theorem 1.3,
$\mu^{{(1)}}_{{1,\operatorname{\mathrm{ex}}}}(\mathfrak{D})=\mu_{1}(\mathfrak{D})$ $\geq\frac{1}{32}\frac{\operatorname{\mathrm{Vol}}(\mathfrak{D}_{1}\cap\mathfrak{D}_{2})}{\operatorname{\mathrm{Vol}}(M)}\min\left\{\mu_{1}(\mathfrak{D}_{1}),\mu_{1}(\mathfrak{D}_{2})\right\}$
$\geq K\left(\frac{R_{c}}{D}\right)^{n}\min\left\{\frac{1}{R_{c}^{2}}\left(\frac{R_{c}}{D}\right)^{3n+2}\left(\frac{R_{c}}{R_{h}}\right)^{7n-5},\frac{1}{R_{c}^{2}}\right\}$
$\geq K\frac{1}{R_{c}^{2}}\left(\frac{R_{c}}{D}\right)^{n}\min\left\{1,\left(\frac{R_{c}}{D}\right)^{3n+2}\left(\frac{R_{c}}{R_{h}}\right)^{7n-5}\right\}.$
Therefore, for $n\geq 2$, we have for any annular region,
$\mu^{{(1)}}_{{1,\operatorname{\mathrm{ex}}}}$ $\geq K\frac{1}{R_{c}^{2}}\min\left\{\left(\frac{R_{c}}{D}\right)^{3n+2}\left(\frac{R_{c}}{R_{h}}\right)^{7n-5},\left(\frac{R_{c}}{D}\right)^{n}\min\left\{1,\left(\frac{R_{c}}{D}\right)^{3n+2}\left(\frac{R_{c}}{R_{h}}\right)^{7n-5}\right\}\right\}$
$\geq K\frac{1}{R_{c}^{2}}\left(\frac{R_{c}}{D}\right)^{n}\min\left\{1,\left(\frac{R_{c}}{D}\right)^{3n+2}\left(\frac{R_{c}}{R_{h}}\right)^{7n-5}\right\},$
thus concluding the proof of Theorem 1.1.∎
Proof of Theorem 1.3: Let $f$ be an eigenfunction associated with ${\mu_{1}}(M)$, normalised by
$\int_{M}f^{2}=1,$
and $\bar{f}_{U}$, $\bar{f}_{V}$ be the averages of $f$ over $U$ and $V$, respectively, i.e.,
$\bar{f}_{U}=\frac{\int_{U}f}{\operatorname{\mathrm{Vol}}(U)}\qquad\text{and}\qquad\bar{f}_{V}=\frac{\int_{V}f}{\operatorname{\mathrm{Vol}}(V)}.$
Then, $f-\bar{f}_{U}$ and $f-\bar{f}_{V}$ are $L^{2}$-orthogonal to the constants, and the variational characterisation for the first positive Neumann eigenvalue gives,
${\mu_{1}}(U)\int_{U\cap V}(f-\bar{f}_{U})^{2}\leq{\mu_{1}}(U)\int_{U}(f-\bar{f}_{U})^{2}\leq\int_{U}|df|^{2}\leq\int_{M}|df|^{2}={\mu_{1}}(M),$ (3.16)
and
${\mu_{1}}(V)\int_{U\cap V}(f-\bar{f}_{V})^{2}\leq{\mu_{1}}(M).$ (3.17)
By writing
$|\bar{f}_{U}-\bar{f}_{V}|=\frac{1}{\operatorname{\mathrm{Vol}}(U\cap V)^{\frac{1}{2}}}\left(\int_{U\cap V}\left((f-\bar{f}_{U})-(f-\bar{f}_{V})\right)^{2}\right)^{\frac{1}{2}},$
we obtain from inequalities (3.16) and (3.17) that
$|\bar{f}_{U}-\bar{f}_{V}|\leq\left(\frac{{\mu_{1}}(M)}{\operatorname{\mathrm{Vol}}(U\cap V)}\right)^{\frac{1}{2}}\left(\frac{1}{\sqrt{{\mu_{1}}(U)}}+\frac{1}{\sqrt{{\mu_{1}}(V)}}\right).$ (3.18)
From (3.17),
$\left|\left(\int_{V}f^{2}\right)^{\frac{1}{2}}-\operatorname{\mathrm{Vol}}(V)^{\frac{1}{2}}|\bar{f}_{V}|\right|\leq\left(\int_{V}(f-\bar{f}_{V})^{2}\right)^{\frac{1}{2}}\leq\sqrt{\frac{{\mu_{1}}(M)}{{\mu_{1}}(V)}}.$
Thus,
$\operatorname{\mathrm{Vol}}(V)^{\frac{1}{2}}|\bar{f}_{V}|\geq\left(\int_{V}f^{2}\right)^{\frac{1}{2}}-\sqrt{\frac{{\mu_{1}}(M)}{{\mu_{1}}(V)}}.$
Without loss of generality, we may assume that
$\int_{V}f^{2}\geq\frac{1}{2}.$
Then,
$|\bar{f}_{V}|\geq\frac{1}{\operatorname{\mathrm{Vol}}(V)^{\frac{1}{2}}}\left(\frac{1}{\sqrt{2}}-\sqrt{\frac{{\mu_{1}}(M)}{{\mu_{1}}(V)}}\right).$
If $\frac{1}{\sqrt{2}}-\sqrt{\frac{{\mu_{1}}(M)}{{\mu_{1}}(V)}}\leq 0$, we obtain
${\mu_{1}}(M)\geq\frac{1}{2}{\mu_{1}}(V),$
which implies the required lower bound of the theorem. Otherwise, since $\operatorname{\mathrm{Vol}}(V)\leq\operatorname{\mathrm{Vol}}(M)$,
$|\bar{f}_{V}|\geq{\frac{1}{\operatorname{\mathrm{Vol}}(M)^{\frac{1}{2}}}}\left(\frac{1}{\sqrt{2}}-\sqrt{\frac{{\mu_{1}}(M)}{{\mu_{1}}(V)}}\right).$ (3.19)
On the other hand, $f$ is orthogonal to constants, so
$0=\int_{M}f=\int_{U}f+\int_{V\setminus U}f=\operatorname{\mathrm{Vol}}(U)(\bar{f}_{U}-\bar{f}_{V})+\operatorname{\mathrm{Vol}}(M)\bar{f}_{V}+\int_{V\setminus U}f-\operatorname{\mathrm{Vol}}(V\setminus U)\bar{f}_{V},$
i.e.,
$0=\operatorname{\mathrm{Vol}}(U)(\bar{f}_{U}-\bar{f}_{V})+\operatorname{\mathrm{Vol}}(M)\bar{f}_{V}+\int_{V\setminus U}(f-\bar{f}_{V}).$
Therefore,
$|\bar{f}_{V}|\leq\frac{1}{\operatorname{\mathrm{Vol}}(M)}\left(\operatorname{\mathrm{Vol}}(U)|\bar{f}_{U}-\bar{f}_{V}|+\int_{V\setminus U}|f-\bar{f}_{V}|\right).$
Using the Cauchy-Schwarz inequality and (3.17), we get
$|\bar{f}_{V}|\leq\frac{1}{\operatorname{\mathrm{Vol}}(M)}\left(\operatorname{\mathrm{Vol}}(U)|\bar{f}_{U}-\bar{f}_{V}|+\operatorname{\mathrm{Vol}}(V\setminus U)^{\frac{1}{2}}\sqrt{\frac{{\mu_{1}}(M)}{{\mu_{1}}(V)}}\right),$
which gives
$|\bar{f}_{V}|\leq|\bar{f}_{U}-\bar{f}_{V}|+{\frac{1}{\sqrt{\operatorname{\mathrm{Vol}}(M)}}}\sqrt{\frac{{\mu_{1}}(M)}{{\mu_{1}}(V)}}.$ (3.20)
Finally, we have from (3.18), (3.19) and (3.20) that
${\frac{1}{\sqrt{\operatorname{\mathrm{Vol}}(M)}}}\left(\frac{1}{\sqrt{2}}-\sqrt{\frac{{\mu_{1}}~(M)}{{\mu_{1}}(V)}}\right)\\ \leq\frac{\sqrt{{\mu_{1}}(M)}}{\sqrt{\operatorname{\mathrm{Vol}}(U\cap V)}}\left(\frac{1}{\sqrt{{\mu_{1}}(U)}}+\frac{1}{\sqrt{{\mu_{1}}(V)}}\right)+{\frac{1}{\sqrt{\operatorname{\mathrm{Vol}}(M)}}}\sqrt{\frac{{\mu_{1}}(M)}{{\mu_{1}}(V)}},$
which implies that
$\frac{1}{\sqrt{2}}$ $\leq 2\left(1+\left(\frac{\operatorname{\mathrm{Vol}}(M)}{\operatorname{\mathrm{Vol}}(U\cap V)}\right)^{\frac{1}{2}}\right)\left(\frac{{\mu_{1}}(M)}{\min\{{\mu_{1}}(U),{\mu_{1}}(V)\}}\right)^{\frac{1}{2}}$
$\leq 4\left(\frac{\operatorname{\mathrm{Vol}}(M)}{\operatorname{\mathrm{Vol}}(U\cap V)}\right)^{\frac{1}{2}}\left(\frac{{\mu_{1}}(M)}{\min\{{\mu_{1}}(U),{\mu_{1}}(V)\}}\right)^{\frac{1}{2}}.$
The theorem follows. ∎
4. Construction of small eigenvalues on annular regions
In this section, we prove Theorem 1.5. In view of the correspondence between the exact part of the $p$-spectrum and the coexact part of the (p1)(p-1)-spectrum, it will suffice to construct, for each $p\in\{0,\dots,n-2\}$, a sequence of domains $(\mathcal{A}_{\epsilon}^{p})_{{\epsilon}\in(0,\frac{1}{2})}$ such that $\lim_{{\epsilon}\to 0}\mu^{(p)}_{1,\operatorname{\mathrm{co}}}(\mathcal{A}_{\epsilon}^{p})=0$. Note that $\mu^{(n-1)}_{1,\operatorname{\mathrm{co}}}$ would be bounded below since
$\mu^{(n-1)}_{1,\operatorname{\mathrm{co}}}=\mu^{{(n)}}_{{1,\operatorname{\mathrm{ex}}}}=\mu^{(n)}_{1}=\lambda_{1},$
and the first Dirichlet eigenvalue $\lambda_{1}$ is uniformly bounded below for domains with bounded diameter, by the Faber-Krahn inequality. The main idea is to create an extra dimension of cohomology on the limiting domain as the contact radius tends to zero. We then use the corresponding harmonic form to construct on the domains appropriate test forms, the supports of whose exterior derivatives have their volumes converging to zero at a required rate, and estimate the associated Rayleigh quotients that appear in the variational characterisation for the first coexact eigenvalue.
Figure 5. $\mathcal{A}^{p}_{\epsilon}$, $\mathcal{A}^{p}_{0}$, and cross-section of $\mathcal{A}^{p}_{\epsilon}$ for $n=3$, $p=1$
Writing $\mathbb{R}^{n}=\mathbb{R}^{n-p-1}\times\mathbb{R}^{p+1}$, we use the cylindrical coordinates given by the combination of polar coordinates $(\rho,{\tilde{\theta}}_{1},\dots,{\tilde{\theta}}_{n-p-2})$ on $\mathbb{R}^{n-p-1}$ with the volume element $\rho^{n-p-2}\,d\rho\,d{\tilde{\theta}}_{1}\dots d{\tilde{\theta}}_{n-p-2}$ and $(r,\theta_{1},\dots,\theta_{p})$ on $\mathbb{R}^{p+1}$ with the volume element $r^{p}\,dr\,d\theta_{1},\dots d\theta_{p}$. For ${\epsilon}\in[0,\frac{1}{2})$, define $\mathcal{A}^{p}_{\epsilon}\subset\mathbb{R}^{n}$ to be the interior of the region given by
[…]
after smoothening it around $\left\{(\rho,r)=\left(0,2\left(\frac{1}{2}+{\epsilon}^{n}\right)^{2}\right)\right\}$, which can be done without affecting it beyond an arbitrarily small neighbourhood and such that the convexity of the domain and the tangency of the boundary to $\theta_{i}$-directions are preserved. In other words, $\mathcal{A}^{p}_{\epsilon}$ is obtained by considering a spherical shell of inner radius $1$ and thickness ${\epsilon}^{n}$, and adding to it the region enclosed by the tangential extensions, at $\{r=\frac{1}{2}\}$, of the parts of its outer boundary. The boundary of $\mathcal{A}^{p}_{\epsilon}$ has $C^{1}$-regularity at the joint involving the tangential extension, while the other joint where the tangential extensions from the two sides meet and close the boundary can be smoothened, as mentioned above. See Figure 5.
Clearly, the diameter of $\mathcal{A}^{p}_{\epsilon}$ and the radius of the hole are bounded independent of $\epsilon$}. Note that $\{\mathcal{A}^{p}_{\epsilon}\}_{{\epsilon}>0}$ are homotopic to $\mathbb{S}^{n-1}$} and have the cohomology
$H^{p}(\mathcal{A}^{p}_{\epsilon})=\begin{cases}0&\text{if }p\neq 0,\\ \mathbb{R}&\text{if }p=0,\end{cases}$
whereas the limiting domain $\mathcal{A}^{p}_{0}$ is homotopic to $\mathbb{S}^{p}$} and hence
$H^{p}(\mathcal{A}^{p}_{0})=\begin{cases}\mathbb{R}&\text{if }p\neq 0,\\ \mathbb{R}^{2}&\text{if }p=0.\end{cases}$
If $p\neq 0$, we may write $\mathbb{R}^{n}=\mathbb{R}^{n-p-1}\times\mathbb{R}^{p+1}$, equipping $\mathbb{R}^{n-p-1}$ with the Cartesian coordinates $(x_{1},\dots,x_{n-p-1})$ and $\mathbb{R}^{p+1}$ with the spherical coordinates $(r,\theta_{1},\dots,\theta_{p})$, where $\theta_{j}\in[0,\pi]$ for $j\in\{1,\dots,p-1\}$ and $\theta_{p}\in[0,2\pi)$. The metric in these coordinates takes the form
$ds^{2}=\bigoplus_{i=1}^{n-p-1}dx_{i}^{2}\oplus dr^{2}\oplus r^{2}\left(\bigoplus_{j=1}^{p}\alpha_{j}^{2}\,d\theta_{j}^{2}\right),$
where $\{\alpha_{j}\}$ are smooth functions of $\{\theta_{j}\}$ given by
$\alpha_{j}(\theta_{1},\dots,\theta_{p})\coloneqq\prod_{k=1}^{j-1}\sin\theta_{k},\qquad j\in\{1,\dots,p\}.$
Then, for $p\neq 0$, the $p$-form
$\omega_{p}\coloneqq\alpha_{1}\dots\alpha_{p}\,d\theta_{1}\wedge\dots\wedge d\theta_{p}$
defined on $\mathbb{R}^{n-p-1}\times(\mathbb{R}^{p+1}\setminus\{0\})$ is a harmonic representative of its $p^{\text{th}}$ cohomology:
$d\omega_{p}=d(\alpha_{1}\dots\alpha_{p})\wedge d\theta_{1}\wedge\dots\wedge d\theta_{p}=0$
and
$\delta\omega_{p}=\pm\star d\star\omega_{p}=\pm\star d(r^{-p}dr\wedge dx_{1}\wedge\dots\wedge dx_{n-p-1})=0.$
The restriction of $\omega_{p}$ to $\mathcal{A}^{p}_{0}$ is a non-zero harmonic form satisfying the absolute boundary conditions, since the $\{\theta_{i}\}$ directions are parallel to the boundary.
If $p=0$, we consider the function $\omega_{0}$ that equals $1$ on one of the connected components and 1-1 on the other.
Let $\chi_{\epsilon}$ be a smooth, non-negative, increasing function defined on $\mathbb{R}_{\geq 0}$ such that it is equal to 0 on $0,\frac{{\epsilon}}{3}]$ and $1$ on $\frac{2{\epsilon}}{3},\infty)$. We construct the family of test forms
$\omega_{p,{\epsilon}}\coloneqq\chi_{\epsilon}(r)\omega_{p}$
defined on $\mathcal{A}^{p}_{\epsilon}$, for ${\epsilon}>0$. Note that the $p$-form $\omega_{p,{\epsilon}}$ is tangential to the boundary of $\mathcal{A}^{p}_{\epsilon}$ (i.e., $i_{\mathpzc{n}}\omega_{p,{\epsilon}}=0$ on $\partial\mathcal{A}^{p}_{\epsilon}$), and when $p=0$, it is orthogonal to the constant functions on $\mathcal{A}^{0}_{\epsilon}$. Furthermore, it is also coclosed:
$\delta\omega_{p,{\epsilon}}=\delta(\chi_{\epsilon}\omega_{p})=\chi_{\epsilon}\,\delta\omega_{p}-i_{\nabla\chi_{\epsilon}}\omega_{p}=0,$
because $\omega_{p}$ is coclosed and $\nabla\chi_{\epsilon}$ is orthogonal to the angular directions corresponding to $\{\theta_{j}\}$. We intend to show that
$\lim_{{\epsilon}\to 0}\frac{\int_{\mathcal{A}^{p}_{\epsilon}}|d\omega_{p,{\epsilon}}|^{2}\,\operatorname{d\hskip-1.00006ptV}}{\int_{\mathcal{A}^{p}_{\epsilon}}|\omega_{p,{\epsilon}}|^{2}\,\operatorname{d\hskip-1.00006ptV}}=0.$
Since $\chi_{\epsilon}=1$ on $\frac{1}{2},\infty)$ for every ${\epsilon}\in(0,\frac{1}{2})$, we have
$\int_{\mathcal{A}^{p}_{\epsilon}}|\omega_{p,{\epsilon}}|^{2}\,\operatorname{d\hskip-1.00006ptV}\geq\int_{\mathcal{A}^{p}_{0}}|\omega_{p}|^{2}\,\operatorname{d\hskip-1.00006ptV}>0,$
i.e., the denominator in the above limit is bounded away from 0, independent of $\epsilon$}. Hence, it would suffice to show that
$\lim_{{\epsilon}\to 0}\int_{\mathcal{A}^{p}_{\epsilon}}|d\omega_{p,{\epsilon}}|^{2}\,\operatorname{d\hskip-1.00006ptV}=0.$
We proceed as follows.
$d\omega_{p,{\epsilon}}=d(\chi_{\epsilon}\omega_{p})=d(\chi_{\epsilon})\wedge\omega_{p}\pm\chi_{\epsilon}\wedge d\omega_{p}=\chi_{\epsilon}^{\prime}(r)\,dr\wedge\omega_{p}.$
So,
$|d\omega_{p,{\epsilon}}|^{2}\,\operatorname{d\hskip-1.00006ptV}$ $=(\chi^{\prime}_{\epsilon})^{2}(dr\wedge\omega_{p})\wedge\star(dr\wedge\omega_{p})$
$=\pm(\chi^{\prime}_{\epsilon})^{2}(dr\wedge\omega_{p})\wedge(r^{-p}dx_{1}\wedge\dots\wedge dx_{n-p-1})$
$=(\chi^{\prime}_{\epsilon})^{2}r^{-2p}\operatorname{d\hskip-1.00006ptV}.$
Choosing the function $\chi$ defined earlier such that $\|\chi^{\prime}\|_{\infty}\leq K{\epsilon}^{-1}$, we get
$\int_{\mathcal{A}^{p}_{\epsilon}}|d\omega_{p,{\epsilon}}|^{2}\,\operatorname{d\hskip-1.00006ptV}$ $\leq K{\epsilon}^{-2p-2}\operatorname{\mathrm{Vol}}\left(\operatorname{Supp}\mathcal{A}^{p}_{\epsilon}\right)$
$=K{\epsilon}^{-2p-2}\operatorname{\mathrm{Vol}}\left(\mathcal{A}^{p}_{\epsilon}\cap\left\{\frac{{\epsilon}}{3}\leq r\leq\frac{2{\epsilon}}{3}\right\}\right).$
The volume term on the RHS is
$\operatorname{\mathrm{Vol}}\left(\mathcal{A}^{p}_{\epsilon}\cap\left\{\frac{{\epsilon}}{3}\leq r\leq\frac{2{\epsilon}}{3}\right\}\right)$ $=K(n,p)\int_{r=\frac{{\epsilon}}{3}}^{\frac{2{\epsilon}}{3}}\int_{\rho=\sqrt{1-r^{2}}}^{\sqrt{(1+{\epsilon}^{n})^{2}-r^{2}}}r^{p}\rho^{n-p-2}\,d\rho\,dr$
$=K\int_{\frac{{\epsilon}}{3}}^{\frac{2{\epsilon}}{3}}r^{p}\left[\left((1+{\epsilon}^{n})^{2}-r^{2}\right)^{\frac{n-p-1}{2}}-\left(1-r^{2}\right)^{\frac{n-p-1}{2}}\right]\,dr,$
which can be estimated as follows. We have that
$r=O\left({\epsilon}\right)\qquad\text{and}\qquad(1+{\epsilon}^{n})^{2}=1+2{\epsilon}^{n}+O\left({\epsilon}^{2n}\right),$
so
$\frac{(1+{\epsilon}^{n})^{2}-1}{1-r^{2}}=O\left({\epsilon}^{n}\right).$
Then
$\left((1+{\epsilon}^{n})^{2}-r^{2}\right)^{\frac{n-p-1}{2}}=(1-r^{2})^{\frac{n-p-1}{2}}\left[1+\frac{1}{2}(n-p-1)\frac{(1+{\epsilon}^{n})^{2}-1}{1-r^{2}}+O\left({\epsilon}^{2n}\right)\right],$
implying that
$\operatorname{\mathrm{Vol}}\left(\operatorname{Supp}\mathcal{A}^{p}_{\epsilon}\right)=\int_{\frac{{\epsilon}}{3}}^{\frac{2{\epsilon}}{3}}r^{p}\,f(r)\,dr,$
with $f(r)=O\left({\epsilon}^{n}\right)$. Hence,
$\operatorname{\mathrm{Vol}}\left(\operatorname{Supp}\mathcal{A}^{p}_{\epsilon}\right)=O\left({\epsilon}^{n+p+1}\right),$
giving that
$\int_{\mathcal{A}^{p}_{\epsilon}}|d\omega_{p,{\epsilon}}|^{2}\,\operatorname{d\hskip-1.00006ptV}=O\left({\epsilon}^{n-p-1}\right).$
Since $p\leq n-2$,
$\lim_{{\epsilon}\to 0}\int_{\mathcal{A}^{p}_{\epsilon}}|d\omega_{p,{\epsilon}}|^{2}\,\operatorname{d\hskip-1.00006ptV}=0.$
Thus,
$\lim_{{\epsilon}\to 0}\frac{\int_{\mathcal{A}^{p}_{\epsilon}}|d\omega_{p,{\epsilon}}|^{2}\,\operatorname{d\hskip-1.00006ptV}}{\int_{\mathcal{A}^{p}_{\epsilon}}|\omega_{p,{\epsilon}}|^{2}\,\operatorname{d\hskip-1.00006ptV}}=0,$
which from the variational characterisation
$\mu^{(p)}_{1,\operatorname{\mathrm{co}}}(M)=\inf_{\begin{subarray}{c}\phi\in\Omega^{p}(M)\\ \delta\phi=0\\ i_{\mathpzc{n}}\phi=0\end{subarray}}\frac{\|d\phi\|^{2}_{L^{2}\Omega^{p+1}(M)}}{\|\phi\|^{2}_{L^{2}\Omega^{p}(M)}}\qquad(\text{for $\phi$ with finite $H^{1}$-norm}),$
implies that $\mu^{(p)}_{1,\operatorname{\mathrm{co}}}(\mathcal{A}^{p}_{\epsilon})\to 0$ as ${\epsilon}\to 0$. ∎
5. Eigenvalue lower bounds for convex domains with multiple holes
In this section, we first present lower bounds for certain higher order exact eigenvalues of convex domains with multiple holes of varying sizes, with no further geometric assumptions on the domain. Next, we prove a generalised version of Theorem 1.6 which gives lower bounds for the first exact eigenvalue under additional geometric hypothesis on the domain. We also construct a family of convex domains with multiples holes, satisfying the hypothesis of Theorem 5.3, having arbitrarily small eigenvalues as the partition parameter tends to zero, with the diameter of the domain and the radius of the largest hole uniformly bounded.
Let $\mathfrak{D}\subset\mathbb{R}^{n}$ be a perforated domain with $\mathfrak{h}$} many holes, that is of the form
$\mathfrak{D}={\widetilde{\mathfrak{D}}}\setminus\bigsqcup_{i=1}^{\mathfrak{h}}\overline{\mathbb{B}_{c_{i},r_{i}}},$
where $\widetilde{\mathfrak{D}}$} is a convex domain in $\mathbb{R}^{n}$ and $\mathbb{B}_{c_{i},r_{i}}$ denotes the Euclidean $n$-ball of radius $r_{i}>0$, centered at $c_{i}\in\mathbb{R}^{n}$. Here, $\{c_{i}\}$ and $\{r_{i}\}$ are assumed to be such that the closures of the balls are mutually disjoint and strictly contained in $\widetilde{\mathfrak{D}}$}.
Figure 6. Laguerre-Voronoi diagram
Consider the Laguerre-Voronoi diagram (also known as the power diagram)  [AKL13, Section 6.2] of $\mathfrak{D}$, with centers $\{c_{i}\}_{i=1}^{\mathfrak{h}}$ and the corresponding weights $\{r_{i}^{2}\}_{i=1}^{\mathfrak{h}}$. This gives a partition of $\mathfrak{D}$ into $\mathfrak{h}$} many subregions (known as “cells”) given by
$\mathcal{C}_{i}\coloneqq\left\{x\in\overline{\mathfrak{D}}:\|x-c_{i}\|^{2}-r_{i}^{2}\leq\|x-c_{j}\|^{2}-r_{j}^{2},\,\,\forall\,j\in\{1,\dots,{\mathfrak{h}}\}\right\},$
such that $\{\mathcal{C}_{i}\cup\mathbb{B}_{c_{i},r_{i}}\}$ are convex. See Figure 6.
We define the partition parameter of $\mathfrak{D}$ associated with the Laguerre-Voronoi diagram to be
$R_{\scriptscriptstyle\hskip-0.73616pt\mathcal{P}}(\mathfrak{D})\coloneqq\min_{i\in\{1,\dots,{\mathfrak{h}}\}}R_{c}(\mathcal{C}_{i}),$
where $R_{c}$ is the contact radius defined earlier for annular domains. Let $k_{p}$ denote the number of non-empty intersections of order $p$ in the Laguerre-Voronoi partition of $\mathfrak{D}$, and, $\hat{r}_{h}$ and $\hat{R}_{h}$ denote the smallest and the largest of the radii of the holes.
Theorem 5.1.
Let $\mathfrak{D}\subset\mathbb{R}^{n}$ be a smooth, bounded, convex domain with $\mathfrak{h}$} many holes, as above. Then the $(1+k_{p})^{\text{th}}$ exact eigenvalue of $\mathfrak{D}$ satisfies the following lower bounds.
  1. (i)
    For $n\geq 3$ and $p\in\{2,\dots,n-1\}$,
    $\mu^{{(p)}}_{{1+k_{p},\operatorname{\mathrm{ex}}}}(\mathfrak{D})\geq\frac{K}{{\mathfrak{h}}^{2p-1}}\frac{1}{D^{2}}\left(\frac{R_{\scriptscriptstyle\hskip-0.73616pt\mathcal{P}}}{D}\right)^{2n^{2}+n-6}\min\left\{1,{\frac{1}{R_{\scriptscriptstyle\hskip-0.73616pt\mathcal{P}}^{2(n-3)}}}\right\}\min\left\{1,\left(\frac{R_{\scriptscriptstyle\hskip-0.73616pt\mathcal{P}}}{\hat{R}_{h}}\right)^{3(n-1)(n+1)}\right\}.$
  2. (ii)
    For $n\geq 2$ and $p=1$,
    $\mu^{{(1)}}_{{1+k_{1},\operatorname{\mathrm{ex}}}}(\mathfrak{D})\geq\frac{K}{{\mathfrak{h}}}\frac{1}{D^{2}}\left(\frac{R_{\scriptscriptstyle\hskip-0.73616pt\mathcal{P}}}{D}\right)^{n-2}\min\left\{1,\left(\frac{R_{\scriptscriptstyle\hskip-0.73616pt\mathcal{P}}}{D}\right)^{3n+2}\left(\frac{R_{\scriptscriptstyle\hskip-0.73616pt\mathcal{P}}}{\hat{R}_{h}}\right)^{7n-5}\right\}.$
Although we use the Laguerre-Voronoi partition in Theorem 5.1 (and Theorem 5.3 below) to have an “optimal” value of the partition parameter, especially when the holes are of identical size, the lower bounds also hold if we consider any partition of $\mathfrak{D}$ into $\mathfrak{h}$} many pieces, each of which is a convex domain with a hole in the interior, and use the partition parameter and $k_{p}$ corresponding to it.
In order to have the best possible partition parameter $R_{\scriptscriptstyle\hskip-0.73616pt\mathcal{P}}$ that we use in our estimates, it may seem more natural to consider the additively weighted Voronoi diagram (whose partition parameter would be equal to the contact radius $\hat{R}_{c}$) where the cells are given by
$\mathfrak{C}_{i}$ $\coloneqq\left\{x\in\overline{\mathfrak{D}}:\operatorname{\mathrm{Dist}}(x,\mathbb{B}_{c_{i},r_{i}})\leq\operatorname{\mathrm{Dist}}(x,\mathbb{B}_{c_{j},r_{j}}),\,\forall\,j\in\{1,\dots,{\mathfrak{h}}\}\right\}$
$=\left\{x\in\overline{\mathfrak{D}}:\|x-c_{i}\|-r_{i}\leq\|x-c_{j}\|-r_{j},\,\forall\,j\in\{1,\dots,{\mathfrak{h}}\}\right\},$
but the outer boundaries of these cells are no more convex, unless the holes are of identical radii. The convexity of the outer boundary is important for our purposes as we intend to use in our proof the lower bounds derived earlier for convex domains minus a hole. Nevertheless, the Laguerre-Voronoi diagram coincides with the standard Voronoi diagram when the holes are of identical size.
Remark 5.2.
The contact radius for convex domains with identical holes defined earlier is also applicable for holes of varying size:
$\hat{R}_{c}(\mathfrak{D})=\min\left\{r_{c},\frac{d_{h}}{2}\right\},$
where $r_{c}$ is the distance between the outer boundary and a hole closest to it and $d_{h}$ is the shortest of the distances between any two holes. Further, if $\hat{d}_{h}$ denotes the largest of the distances between any two holes, an elementary computation shows that the partition parameter satisfies
$\left(\frac{d_{h}+2\hat{r}_{h}}{\hat{d}_{h}+2\hat{R}_{h}}\right)\hat{R}_{c}\leq R_{\scriptscriptstyle\hskip-0.73616pt\mathcal{P}}\leq\hat{R}_{c},$
and that $R_{\scriptscriptstyle\hskip-0.73616pt\mathcal{P}}=\hat{R}_{c}$ when the holes are of identical size.
These lower bounds from Theorem 5.1 in general do not give lower bounds for the smallest exact eigenvalue $\mu^{{(p)}}_{{1,\operatorname{\mathrm{ex}}}}$, unless the Laguerre-Voronoi partition of $\mathfrak{D}$ (or, if exists, any other partition with the cells having a convex outer boundary) has no $p$-intersections. However, we can derive the lower bounds of Theorem 5.3 below for $\mu^{{(p)}}_{{1,\operatorname{\mathrm{ex}}}}$ under additional geometric hypothesis on $\mathfrak{D}$, using an induction-like approach.
Hypothesis (H).
The Laguerre-Voronoi diagram $\{\mathcal{C}_{i}\}_{i=1}^{\mathfrak{h}}$ of the domain $\mathfrak{D}$, up to reordering of the cells, is such that
$\bigcup_{i=1}^{\ell}\,\mathcal{C}_{i}\cup\,\mathbb{B}_{c_{i},r_{i}}$
is convex for each $\ell\in\{1,\dots,{\mathfrak{h}}\}$.
Theorem 5.3.
Let $\mathfrak{D}\subset\mathbb{R}^{n}$ be smooth, bounded, convex domain with $\mathfrak{h}$} many holes, as above. Suppose that $\mathfrak{D}$ satisfies the geometric condition given by Hypothesis (H). Then we have the following lower bounds for the first exact eigenvalue.
  1. (i)
    For $n\geq 3$ and $p\in\{2,\dots,n-1\}$,
    $\mu^{{(p)}}_{{1,\operatorname{\mathrm{ex}}}}(\mathfrak{D})\geq K\frac{1}{D^{2}}\left(\frac{R_{\scriptscriptstyle\hskip-0.73616pt\mathcal{P}}}{D}\right)^{2n^{2}-n-4+2{\mathfrak{h}}}\min\left\{1,\left(\frac{R_{\scriptscriptstyle\hskip-0.73616pt\mathcal{P}}}{\hat{R}_{h}}\right)^{3(n-1)(n+1)}\right\}.$
  2. (ii)
    For $n\geq 2$ and $p=1$,
    $\mu^{{(1)}}_{{1,\operatorname{\mathrm{ex}}}}(\mathfrak{D})\geq K\frac{1}{D^{2}}\left(\frac{R_{\scriptscriptstyle\hskip-0.73616pt\mathcal{P}}\hat{r}_{h}^{n-1}}{D^{n}}\right)^{{\mathfrak{h}}-1}\left(\frac{R_{\scriptscriptstyle\hskip-0.73616pt\mathcal{P}}}{D}\right)^{n-2}\min\left\{1,\left(\frac{R_{\scriptscriptstyle\hskip-0.73616pt\mathcal{P}}}{D}\right)^{3n+2}\left(\frac{R_{\scriptscriptstyle\hskip-0.73616pt\mathcal{P}}}{\hat{R}_{h}}\right)^{7n-5}\right\}.$
The exponents of the ratios $R_{\scriptscriptstyle\hskip-0.73616pt\mathcal{P}}/D$ and $R_{\scriptscriptstyle\hskip-0.73616pt\mathcal{P}}\hat{r}_{h}^{n-1}/D^{n}$ in Theorem 5.3 depend on the number of holes, in contrast to the lower bounds of Theorem 5.1.
Remark 5.4.
Note that Hypothesis (H) is always true when $\mathfrak{D}$ has at most two holes. Furthermore, Theorem 5.3 can be adapted to give similar lower bounds for a larger class of domains than those satisfying this hypothesis, for instance, when the cells of the Laguerre-Voronoi partition may be arranged into a finite nested sequence, with the subset of the cells from each inner sequence satisfying Hypothesis (H) and the set of unions of the cells in each inner sequence themselves satisfying an analogue of Hypothesis (H).
The following theorem emphasises the role of the partition parameter $R_{\scriptscriptstyle\hskip-0.73616pt\mathcal{P}}$ in the lower bounds of Theorem 5.3, for $n\geq 3$ and $p\in\{2,\dots,n-1\}$.
Theorem 5.5.
Let $n\geq 2$ and $p\in\{1,\dots,n-1\}$. Given $\mathfrak{h}$$\in\mathbb{N}$, there exists a family of Euclidean $C^{1}$-domains with convex outer boundary and $\mathfrak{h}$} many holes, satisfying Hypothesis (H) and having arbitrarily small exact eigenvalue, with the diameter of the domain and the radius of the largest hole uniformly bounded, and the partition parameter tending to zero.
We proceed with the proofs of the above theorems.
Proof of Theorem 5.1. The cells $\{\mathcal{C}_{i}\}$ from the Laguerre-Voronoi partition have trivial cohomologies for $p\in\{1,\dots,n-2\}$. Let $\{{\widetilde{\mathcal{C}}_{i}}\}$ be smooth extensions of $\{\mathcal{C}_{i}\}$ around the parts of their outer boundaries in the interior of $\mathfrak{D}$, small enough not to intersect the hole from any neighbouring cell, such that $\{{\widetilde{\mathcal{C}}_{i}}\cup\mathbb{B}_{c_{i},r_{i}}\}$ are still convex and contained in $\widetilde{\mathfrak{D}}$}.
Observe that the diameters,
$\operatorname{\mathrm{Diam}}\left({\widetilde{\mathcal{C}}_{i}}\right)\leq\operatorname{\mathrm{Diam}}(\mathfrak{D})=D,$
and the contact radii
$R_{c}\left({\widetilde{\mathcal{C}}_{i}}\right)\geq R_{c}(\mathcal{C}_{i})\geq R_{\scriptscriptstyle\hskip-0.73616pt\mathcal{P}}.$
Then we have from Theorem 1.1 that, for $n\geq 3$ and $p\in\{2,\dots n-1\}$,
$\mu^{{(p)}}_{{1,\operatorname{\mathrm{ex}}}}\left({\widetilde{\mathcal{C}}_{i}}\right)$ $\geq K\frac{1}{\operatorname{\mathrm{Diam}}\left({\widetilde{\mathcal{C}}_{i}}\right)^{2}}\left(\frac{R_{c}\left({\widetilde{\mathcal{C}}_{i}}\right)}{\operatorname{\mathrm{Diam}}\left({\widetilde{\mathcal{C}}_{i}}\right)}\right)^{2n^{2}-n-2}\min\left\{1,\left(\frac{R_{c}\left({\widetilde{\mathcal{C}}_{i}}\right)}{r_{i}}\right)^{3(n-1)(n+1)}\right\}$
$\geq K\frac{1}{D^{2}}\left(\frac{R_{\scriptscriptstyle\hskip-0.73616pt\mathcal{P}}}{D}\right)^{2n^{2}-n-2}\min\left\{1,\left(\frac{R_{\scriptscriptstyle\hskip-0.73616pt\mathcal{P}}}{\hat{R}_{h}}\right)^{3(n-1)(n+1)}\right\},\qquad\forall\,i\in\{1,\dots,{\mathfrak{h}}\}.$ (5.1)
Similarly, for $n\geq 2$ and $p=1$,
$\mu^{{(1)}}_{{1,\operatorname{\mathrm{ex}}}}\left({\widetilde{\mathcal{C}}_{i}}\right)\geq K\frac{1}{D^{2}}\left(\frac{R_{\scriptscriptstyle\hskip-0.73616pt\mathcal{P}}}{D}\right)^{n-2}\min\left\{1,\left(\frac{R_{\scriptscriptstyle\hskip-0.73616pt\mathcal{P}}}{D}\right)^{3n+2}\left(\frac{R_{\scriptscriptstyle\hskip-0.73616pt\mathcal{P}}}{\hat{R}_{h}}\right)^{7n-5}\right\},\qquad\forall\,i\in\{1,\dots,{\mathfrak{h}}\}.$ (5.2)
Since $\{{\widetilde{\mathcal{C}}_{i}}\cup\mathbb{B}_{c_{i},r_{i}}\}$ are convex, so is any non-trivial intersection of a subset of them. From the known lower bound for convex domains [Gue04b], we have for any non-trivial intersection,
$\mu^{{(p)}}_{{1,\operatorname{\mathrm{ex}}}}\left({\widetilde{\mathcal{C}}_{i_{1}}}\cap\dots\cap{\widetilde{\mathcal{C}}_{i_{\ell}}}\right)\geq\frac{K}{\operatorname{\mathrm{Diam}}\left({\widetilde{\mathcal{C}}_{i_{1}}}\cap\dots\cap{\widetilde{\mathcal{C}}_{i_{\ell}}}\right)^{2}}\geq\frac{K}{D^{2}},\qquad\forall\,p\in\{0,\dots,n\}.$ (5.3)
Below, we present certain variants of the Cheeger–McGowan lemma [McG93, Lemma 2.3] and [GP95, Lemma 1], which we use in this proof in combination with the above lower bounds for the individual cells.
Lemma 5.6.
Let $(M^{n},g)$ be a compact Riemannian manifold with boundary and $\{U_{i}\}_{i=1}^{k_{0}}$ be an open cover with a partition of unity $\{\rho_{i}\}_{i=1}^{k_{0}}$ subordinate to it, with
$c_{\rho}\coloneqq\max_{i\in\{1,\dots,k_{0}\}}\|\nabla\rho_{i}\|^{2}_{\infty}.$
[…]
  1. (1)
    For $n\geq 2$ and $p=1$,
    $\mu^{{(1)}}_{{1+k_{1},\operatorname{\mathrm{ex}}}}(M)\geq\frac{1}{\sum_{i=1}^{k_{0}}{\frac{1}{\mu^{{(1)}}_{{1,\operatorname{\mathrm{ex}}}}(U_{i})}}}.$
  2. (2)
    Let $n\geq 3$ and $p=2$. Suppose that $H^{1}(U_{ij},\mathbb{R})=0$ for every non-trivial $1$-intersection $U_{ij}$. Then
    $\mu^{{(2)}}_{{1+k_{2},\operatorname{\mathrm{ex}}}}(M)\geq\frac{(8k_{0})^{-1}}{\sum_{i=1}^{k_{0}}\Biggl\{{\frac{1}{\mu^{{(2)}}_{{1,\operatorname{\mathrm{ex}}}}(U_{i})}}+\sum_{j=1}^{k_{0}}\Biggl[\left(\frac{c_{\rho}}{\mu^{{(1)}}_{{1,\operatorname{\mathrm{ex}}}}(U_{ij})}+1\right)\left({\frac{1}{\mu^{{(2)}}_{{1,\operatorname{\mathrm{ex}}}}(U_{i})}}+{\frac{1}{\mu^{{(2)}}_{{1,\operatorname{\mathrm{ex}}}}(U_{j})}}\right)\Biggr]\Biggr\}}.$
  3. (3)
    Let $n\geq 4$ and $p=3$. Suppose that $H^{2}(U_{ij},\mathbb{R})=0$ for every non-trivial $1$-intersection $U_{ij}$ and $H^{1}(U_{ijk},\mathbb{R})=0$ for every non-trivial intersection $2$-intersection $U_{ijk}$. Then
    $\mu^{{(3)}}_{{1+k_{3},\operatorname{\mathrm{ex}}}}(M)\geq\frac{(18k_{0}^{2})^{-1}}{\aleph},$
    where the denominator $\aleph$ is
    $\hskip 28.90036pt\sum_{i=1}^{k_{0}}$ $\Biggl\{{\frac{1}{\mu^{{(3)}}_{{1,\operatorname{\mathrm{ex}}}}(U_{i})}}+\sum_{j=1}^{k_{0}}\Biggl[\left(\frac{c_{\rho}}{\mu^{{(2)}}_{{1,\operatorname{\mathrm{ex}}}}(U_{ij})}+1\right)\left({\frac{1}{\mu^{{(3)}}_{{1,\operatorname{\mathrm{ex}}}}(U_{i})}}+{\frac{1}{\mu^{{(3)}}_{{1,\operatorname{\mathrm{ex}}}}(U_{j})}}\right)$
    $+\sum_{k=1}^{k_{0}}c_{\rho}\left(\frac{c_{\rho}}{\mu^{{(1)}}_{{1,\operatorname{\mathrm{ex}}}}(U_{ijk})}+1\right)\biggl\{\frac{{\frac{1}{\mu^{{(3)}}_{{1,\operatorname{\mathrm{ex}}}}(U_{i})}}+{\frac{1}{\mu^{{(3)}}_{{1,\operatorname{\mathrm{ex}}}}(U_{j})}}}{\mu^{{(2)}}_{{1,\operatorname{\mathrm{ex}}}}(U_{ij})}+\frac{{\frac{1}{\mu^{{(3)}}_{{1,\operatorname{\mathrm{ex}}}}(U_{j})}}+{\frac{1}{\mu^{{(3)}}_{{1,\operatorname{\mathrm{ex}}}}(U_{k})}}}{\mu^{{(2)}}_{{1,\operatorname{\mathrm{ex}}}}(U_{jk})}+\frac{{\frac{1}{\mu^{{(3)}}_{{1,\operatorname{\mathrm{ex}}}}(U_{i})}}+{\frac{1}{\mu^{{(3)}}_{{1,\operatorname{\mathrm{ex}}}}(U_{k})}}}{\mu^{{(2)}}_{{1,\operatorname{\mathrm{ex}}}}(U_{ik})}\biggr\}\Biggr]\Biggr\}.$
Part $(2)$ of this lemma for the case $n\geq 3$, $p=2$ is a direct consequence of [McG93, Lemma 2.3], where it was originally stated for closed manifolds but the proof also holds for compact manifolds with boundary. Part $(1)$ for $n\geq 2$, $p=1$ is a straight forward adaptation of McGowan’s method and is known in the literature, although there seems to be no mention of this lower bound explicitly. The expression in the lower bound of Part $(3)$ for $n\geq 4$, $p=3$ is a new result that we derive adapting McGowan’s proof, motivated by the concluding remark of [McG93, Section 2] that it should be possible to generalise their lemma to higher degree forms. The proof of this case is somewhat technical and is provided in the Appendix.
Remark 5.7.
Using similar arguments for higher degrees $p$, one can derive lower bounds on $\mu^{{(p)}}_{{1+k_{p},\operatorname{\mathrm{ex}}}}(M)$ with more sophisticated expressions, supposing that the $q^{\text{th}}$ cohomology of every non-trivial (pq)(p-q)-intersection is trivial for all $q\in\{1,\dots,p-1\}$. Specifically, the expression for a lower bound on $\mu^{{(p)}}_{{1+k_{p},\operatorname{\mathrm{ex}}}}(M)$ will have a nested combinations of the inverses of $\mu^{{(p-q)}}_{{1,\operatorname{\mathrm{ex}}}}(U_{i_{0}\dots i_{q}})$ for $q\in\{0,\dots,p-1\}$, with a dependence of $k_{0}^{1-p}$ on the number of elements in the open cover of $M$, and the dependence on the partition of unity reflected in the factors
$\left(\frac{c_{\rho}}{\mu^{{(p-1)}}_{{1,\operatorname{\mathrm{ex}}}}(U_{i_{0}i_{1}})}+1\right),\,c_{\rho}\left(\frac{c_{\rho}}{\mu^{{(p-2)}}_{{1,\operatorname{\mathrm{ex}}}}(U_{i_{0}i_{1}i_{2}})}+1\right),\dots,\,c_{\rho}^{p-2}\left(\frac{c_{\rho}}{\mu^{{(p-3)}}_{{1,\operatorname{\mathrm{ex}}}}(U_{i_{0}i_{1}i_{2}i_{3}})}+1\right)$
in the denominator.
Returning to the main proof of Theorem 5.1, consider the open cover $\left\{\operatorname{Int}{\widetilde{\mathcal{C}}_{i}}\right\}_{i=1}^{\mathfrak{h}}$ of $\mathfrak{D}$ and a partition of unity $\{\rho_{i}\}_{i=1}^{{\mathfrak{h}}}$ subordinate to it such that
$c_{\rho}\coloneqq\max_{i\in\{1,\dots,{\mathfrak{h}}\}}\|\nabla\rho_{i}\|^{2}_{\infty}\leq\frac{K}{R_{\scriptscriptstyle\hskip-0.73616pt\mathcal{P}}^{2}}.$
For $n\geq 2$ and $p=1$, it follows from part $(1)$ of Lemma 5.6 and the lower bound (5.2) that
$\mu^{{(1)}}_{{1+k_{1},\operatorname{\mathrm{ex}}}}(\mathfrak{D})\geq\frac{K}{{\mathfrak{h}}}\frac{1}{D^{2}}\left(\frac{R_{\scriptscriptstyle\hskip-0.73616pt\mathcal{P}}}{D}\right)^{n-2}\min\left\{1,\left(\frac{R_{\scriptscriptstyle\hskip-0.73616pt\mathcal{P}}}{D}\right)^{3n+2}\left(\frac{R_{\scriptscriptstyle\hskip-0.73616pt\mathcal{P}}}{\hat{R}_{h}}\right)^{7n-5}\right\}.$
For $n\geq 3$ and $p=2$, we use part $(2)$ of Lemma 5.6, the lower bounds (5.1) and (5.3), and that $c_{\rho}\leq KR_{\scriptscriptstyle\hskip-0.73616pt\mathcal{P}}^{-2}$, to get
$\mu^{{(2)}}_{{1+k_{2},\operatorname{\mathrm{ex}}}}(\mathfrak{D})\geq\frac{K}{{\mathfrak{h}}^{3}}\frac{1}{D^{2}}\left(\frac{R_{\scriptscriptstyle\hskip-0.73616pt\mathcal{P}}}{D}\right)^{2n^{2}-n}\min\left\{1,\left(\frac{R_{\scriptscriptstyle\hskip-0.73616pt\mathcal{P}}}{\hat{R}_{h}}\right)^{3(n-1)(n+1)}\right\}.$
Likewise, for $n\geq 3$ and $p=3$, we use part $(3)$ of Lemma 5.6 to get
$\mu^{{(3)}}_{{1+k_{3},\operatorname{\mathrm{ex}}}}(\mathfrak{D})\geq\frac{K}{{\mathfrak{h}}^{5}}\frac{1}{D^{2}}\left(\frac{R_{\scriptscriptstyle\hskip-0.73616pt\mathcal{P}}}{D}\right)^{2n^{2}-n+2}\min\left\{1,\left(\frac{R_{\scriptscriptstyle\hskip-0.73616pt\mathcal{P}}}{\hat{R}_{h}}\right)^{3(n-1)(n+1)}\right\}.$
For $n\geq 3$ and $p\in\{2,\dots,n-1\}$, in general, we use expression for higher degree forms described in Remark 5.7. Using the lower bounds (5.1) and (5.3), and that
$c_{\rho}^{p-2}\leq K\max\{1,(R_{\scriptscriptstyle\hskip-0.73616pt\mathcal{P}}^{-2})^{p-2}\},$
the denominator in the expression can be bounded above by
$K\max\left\{1,(R_{\scriptscriptstyle\hskip-0.73616pt\mathcal{P}}^{-2})^{p-2}\right\}\left(\frac{R_{\scriptscriptstyle\hskip-0.73616pt\mathcal{P}}^{-2}}{D^{-2}}+1\right){\mathfrak{h}}^{p}\left({\frac{1}{D^{-2}}}\right)^{p-2}\left(\frac{1}{D^{2}}\left(\frac{R_{\scriptscriptstyle\hskip-0.73616pt\mathcal{P}}}{D}\right)^{2n^{2}-n-2}\min\left\{1,\left(\frac{R_{\scriptscriptstyle\hskip-0.73616pt\mathcal{P}}}{\hat{R}_{h}}\right)^{3(n-1)(n+1)}\right\}\right)^{-1}$
which yields
$\mu^{{(p)}}_{{1+k_{p},\operatorname{\mathrm{ex}}}}(\mathfrak{D})$ $\geq\frac{K}{{\mathfrak{h}}^{2p-1}}\frac{R_{\scriptscriptstyle\hskip-0.73616pt\mathcal{P}}^{2}\min\left\{1,R_{\scriptscriptstyle\hskip-0.73616pt\mathcal{P}}^{2(p-2)}\right\}}{D^{2(p-1)}}\frac{1}{D^{2}}\left(\frac{R_{\scriptscriptstyle\hskip-0.73616pt\mathcal{P}}}{D}\right)^{2n^{2}-n-2}\min\left\{1,\left(\frac{R_{\scriptscriptstyle\hskip-0.73616pt\mathcal{P}}}{\hat{R}_{h}}\right)^{3(n-1)(n+1)}\right\}$
$\geq\frac{K}{{\mathfrak{h}}^{2p-1}}\frac{1}{D^{2}}\left(\frac{R_{\scriptscriptstyle\hskip-0.73616pt\mathcal{P}}}{D}\right)^{2n^{2}-n-2+2(p-1)}\min\left\{1,{\frac{1}{R_{\scriptscriptstyle\hskip-0.73616pt\mathcal{P}}^{2(p-2)}}}\right\}\min\left\{1,\left(\frac{R_{\scriptscriptstyle\hskip-0.73616pt\mathcal{P}}}{\hat{R}_{h}}\right)^{3(n-1)(n+1)}\right\}$
$\geq\frac{K}{{\mathfrak{h}}^{2p-1}}\frac{1}{D^{2}}\left(\frac{R_{\scriptscriptstyle\hskip-0.73616pt\mathcal{P}}}{D}\right)^{2n^{2}+n-6}\min\left\{1,{\frac{1}{R_{\scriptscriptstyle\hskip-0.73616pt\mathcal{P}}^{2(n-3)}}}\right\}\min\left\{1,\left(\frac{R_{\scriptscriptstyle\hskip-0.73616pt\mathcal{P}}}{\hat{R}_{h}}\right)^{3(n-1)(n+1)}\right\}.$
Proof of Theorem 5.3. As before, consider the open cover $\{\operatorname{Int}{\widetilde{\mathcal{C}}_{i}}\}_{i=1}^{\mathfrak{h}}$ of $\mathfrak{D}$, where $\{\mathcal{C}_{i}\}_{i=1}^{\mathfrak{h}}$ are ordered as given by Hypothesis (H), and such that $\bigcup_{i=1}^{\ell}\,{\widetilde{\mathcal{C}}_{i}}\cup\,\mathbb{B}_{c_{i},r_{i}}$ is convex for each $\ell\in\{1,\dots,{\mathfrak{h}}\}$. Let $\{\rho_{i}\}_{i=1}^{{\mathfrak{h}}}$ be a partition of unity subordinate to this open cover, satisfying
$c_{\rho}\coloneqq\max_{i\in\{1,\dots,{\mathfrak{h}}\}}\|\nabla\rho_{i}\|^{2}_{\infty}\leq\frac{K}{R_{\scriptscriptstyle\hskip-0.73616pt\mathcal{P}}^{2}}.$
When $n\geq 3$ and $p\in\{2,\dots,n-1\}$, we proceed to prove the following using induction: For each $\ell\in\{1,\dots,{\mathfrak{h}}\}$,
$\mu^{{(p)}}_{{1,\operatorname{\mathrm{ex}}}}\left({\widetilde{\mathcal{C}}_{1}}\cup\dots\cup{\widetilde{\mathcal{C}}_{\ell}}\right)\geq K\frac{1}{D^{2}}\left(\frac{R_{\scriptscriptstyle\hskip-0.73616pt\mathcal{P}}}{D}\right)^{2n^{2}-n-2+2(\ell-1)}\min\left\{1,\left(\frac{R_{\scriptscriptstyle\hskip-0.73616pt\mathcal{P}}}{\hat{R}_{h}}\right)^{3(n-1)(n+1)}\right\}.$
The above statement is true for $\ell=1$ from the lower bound (5.1). Suppose that the induction hypothesis holds for some $\ell\in\{1,\dots,{\mathfrak{h}}-1\}$. Using Lemma 3.7 for the open cover of $\{{\widetilde{\mathcal{C}}_{1}}\cup\dots\cup{\widetilde{\mathcal{C}}_{\ell+1}}\}$ given by $\{{\widetilde{\mathcal{C}}_{1}}\cup\dots\cup{\widetilde{\mathcal{C}}_{\ell}},{\widetilde{\mathcal{C}}_{\ell+1}}\}$, and that $c_{\rho}\leq KR_{c}^{-2}$, we get
$\mu^{{(p)}}_{{1,\operatorname{\mathrm{ex}}}}\left({\widetilde{\mathcal{C}}_{1}}\cup\dots\cup{\widetilde{\mathcal{C}}_{\ell+1}}\right)\geq K\frac{1}{D^{2}}\left(\frac{R_{\scriptscriptstyle\hskip-0.73616pt\mathcal{P}}}{D}\right)^{2n^{2}-n-2+2\ell}\min\left\{1,\left(\frac{R_{\scriptscriptstyle\hskip-0.73616pt\mathcal{P}}}{\hat{R}_{h}}\right)^{3(n-1)(n+1)}\right\},$
hence the induction statement is true for $\ell+1$. The first part of Theorem 5.3 follows.
When $n\geq 2$ and $p=1$, we proceed as follows, again using induction: For each $\ell\in\{1,\dots,{\mathfrak{h}}\}$,
$\mu^{{(1)}}_{{1,\operatorname{\mathrm{ex}}}}\left({\widetilde{\mathcal{C}}_{1}}\cup\dots\cup{\widetilde{\mathcal{C}}_{\ell}}\right)\geq K\frac{1}{D^{2}}\left(\frac{R_{\scriptscriptstyle\hskip-0.73616pt\mathcal{P}}\hat{r}_{h}^{n-1}}{D^{n}}\right)^{\ell-1}\left(\frac{R_{\scriptscriptstyle\hskip-0.73616pt\mathcal{P}}}{D}\right)^{n-2}\min\left\{1,\left(\frac{R_{\scriptscriptstyle\hskip-0.73616pt\mathcal{P}}}{D}\right)^{3n+2}\left(\frac{R_{\scriptscriptstyle\hskip-0.73616pt\mathcal{P}}}{\hat{R}_{h}}\right)^{7n-5}\right\}.$
The above statement is true for $\ell=1$ from the lower bound (5.2). Suppose that the induction hypothesis holds for some $\ell\in\{1,\dots,{\mathfrak{h}}-1\}$. Using Theorem 1.3,
$\mu^{{(1)}}_{{1,\operatorname{\mathrm{ex}}}}\left({\widetilde{\mathcal{C}}_{1}}\cup\dots\cup{\widetilde{\mathcal{C}}_{\ell+1}}\right)\geq\frac{1}{32}\frac{\operatorname{\mathrm{Vol}}\left(({\widetilde{\mathcal{C}}_{1}}\cup\dots\cup{\widetilde{\mathcal{C}}_{\ell}})\cap{\widetilde{\mathcal{C}}_{\ell+1}}\right)}{\operatorname{\mathrm{Vol}}\left({\widetilde{\mathcal{C}}_{1}}\cup\dots\cup{\widetilde{\mathcal{C}}_{\ell+1}}\right)}\min\left\{\mu^{{(1)}}_{{1,\operatorname{\mathrm{ex}}}}\left({\widetilde{\mathcal{C}}_{1}}\cup\dots\cup{\widetilde{\mathcal{C}}_{\ell}}\right),\mu^{{(1)}}_{{1,\operatorname{\mathrm{ex}}}}\left({\widetilde{\mathcal{C}}_{\ell+1}}\right)\right\}.$
We have that
$\operatorname{\mathrm{Vol}}\left({\widetilde{\mathcal{C}}_{1}}\cup\dots\cup{\widetilde{\mathcal{C}}_{\ell+1}}\right)\leq KD^{n},$
while the volume term in the numerator may be estimated as
$\operatorname{\mathrm{Vol}}\left(({\widetilde{\mathcal{C}}_{1}}\cup\dots\cup{\widetilde{\mathcal{C}}_{\ell}})\cap{\widetilde{\mathcal{C}}_{\ell+1}}\right)\geq KR_{\scriptscriptstyle\hskip-0.73616pt\mathcal{P}}r_{\ell+1}^{n-1}\geq KR_{\scriptscriptstyle\hskip-0.73616pt\mathcal{P}}\hat{r}_{h}^{n-1},$
since the extensions ${\widetilde{\mathcal{C}}_{1}},\dots,{\widetilde{\mathcal{C}}_{\ell}}$ may be chosen large enough such that their union contains at least the subset of $\mathcal{C}_{\ell+1}$ given by the cylinder aligned along the perpendicular from $c_{\ell+1}$ to the hyperplane of partition, of radius equal to $r_{\ell+1}$ and height $R_{\scriptscriptstyle\hskip-0.73616pt\mathcal{P}}$, with its centre at a distance of $r_{\ell+1}+\frac{R_{\scriptscriptstyle\hskip-0.73616pt\mathcal{P}}}{2}$ from $c_{\ell+1}$. Using these volume estimates along with the induction hypothesis and the lower bound (5.2), we get
$\mu^{{(1)}}_{{1,\operatorname{\mathrm{ex}}}}\left({\widetilde{\mathcal{C}}_{1}}\cup\dots\cup{\widetilde{\mathcal{C}}_{\ell+1}}\right)\geq K\frac{1}{D^{2}}\left(\frac{R_{\scriptscriptstyle\hskip-0.73616pt\mathcal{P}}\hat{r}_{h}^{n-1}}{D^{n}}\right)^{\ell}\left(\frac{R_{\scriptscriptstyle\hskip-0.73616pt\mathcal{P}}}{D}\right)^{n-2}\min\left\{1,\left(\frac{R_{\scriptscriptstyle\hskip-0.73616pt\mathcal{P}}}{D}\right)^{3n+2}\left(\frac{R_{\scriptscriptstyle\hskip-0.73616pt\mathcal{P}}}{\hat{R}_{h}}\right)^{7n-5}\right\},$
hence the induction statement is true for $\ell+1$. The second part of Theorem 5.3 follows. ∎
Proof of Theorem 5.5. Fix $n\geq 2$, $p\in\{1,\dots,n-1\}$, and $\mathfrak{h}$$\in\mathbb{N}$. Consider the family of domains $\{\mathcal{A}^{p-1}_{\epsilon}\}_{{\epsilon}\in[0,\frac{1}{2})}$ as in the proof of Theorem 1.5 in Section 4, and the cartesian coordinates $(x_{1},\dots,x_{n})$ of $\mathbb{R}^{n}=\mathbb{R}^{n-p}\times\mathbb{R}^{p}$, where $\mathbb{R}^{n-p}$ and $\mathbb{R}^{p}$ are as before. Recall that Section 4 approaches the proof of Theorem 1.5 using the corresponding coexact eigenvalues for convenience, whereas we now directly analyse the exact eigenvalues of degree $p$, hence the choice $\{\mathcal{A}^{p-1}_{\epsilon}\}$ associated with $p$. For sufficiently small $r_{0}>0$ depending on $\mathfrak{h}$}, construct the family of domains $\{\mathcal{A}^{p-1}_{{\epsilon},r}\}$ from $\{\mathcal{A}^{p-1}_{\epsilon}\}$ by removing ${\mathfrak{h}}-1$ non-overlapping balls of radius $r\in(0,r_{0})$ contained in the interior of $\mathcal{A}^{p-1}_{0}$ and centered on the $x_{1}$-axis.
Figure 7.
Clearly, the domains $\{\mathcal{A}^{p-1}_{{\epsilon},r}\}$ satisfy Hypothesis (H), and have uniformly bounded diameter and radius of the largest hole, with the partition parameter tending to zero as ${\epsilon}\to 0$. It then follows from Proposition 3.6 and Theorem 1.5 that
$\lim_{{\epsilon}\to 0}\lim_{r\to 0}\mu^{{(p)}}_{{1,\operatorname{\mathrm{ex}}}}(\mathcal{A}^{p-1}_{{\epsilon},r})=\lim_{{\epsilon}\to 0}\mu^{{(p)}}_{{1,\operatorname{\mathrm{ex}}}}(\mathcal{A}^{p-1}_{\epsilon})=\lim_{{\epsilon}\to 0}\mu^{(p-1)}_{1,\operatorname{\mathrm{co}}}(\mathcal{A}^{p-1}_{\epsilon})=0,$
thereby giving rise to arbitrarily small exact eigenvalues. ∎
Appendix: Generalisation of McGowan’s lemma
We give a proof of part $(3)$ of Lemma 5.6, adapting that of [McG93, Lemma 2.3] to the case of $n\geq 4$, $p=3$. We begin by recalling a variational characterisation for the exact eigenvalues with absolute boundary conditions [Dod82, Proposition 3.1] (see [McG93, Proposition 2.1] for a proof): For $p\in\{1,\dots,n\}$,
$\mu^{{(p)}}_{{k,\operatorname{\mathrm{ex}}}}(M)=\inf_{V_{k}}\sup_{\eta\in V_{k}\setminus\{0\}}\,\sup_{\theta\,:\,\eta=d\theta}\frac{\|\eta\|^{2}_{L^{2}\Omega^{p}(M)}}{\|\theta\|^{2}_{L^{2}\Omega^{p-1}(M)}},$ (5.4)
where $V_{k}$ ranges over all $k$-dimensional subspaces of the space of $L^{2}$-integrable exact $p$-forms on $M$. Let $\mathcal{E}_{k}$ denote the span of the eigenforms corresponding to $\left\{\mu^{{(3)}}_{{1,\operatorname{\mathrm{ex}}}},\dots,\mu^{{(3)}}_{{k,\operatorname{\mathrm{ex}}}}\right\}$. Then for any $\phi\in\mathcal{E}_{1+k_{3}}$ and $\psi$ such that $d\psi=\phi$,
$\mu^{{(3)}}_{{1+k_{3},\operatorname{\mathrm{ex}}}}\geq\frac{\|\phi\|^{2}_{L^{2}\Omega^{3}(M)}}{\|\psi\|^{2}_{L^{2}\Omega^{2}(M)}}.$
So, it will suffice to construct such a pair with the above quotient of the (squares of the) $L^{2}$-norms bounded below by appropriate terms. Given a $\phi$, we can locally construct $\psi$ using the eigenvalues of the pieces $\{U_{i}\}$ and their intersections. We then use a partition of unity for the local-to-global transition employing the Cěch–de Rham formalism, which enforces a choice of $\phi$ to be made in the proof, as we shall see.
[…]
Figure 8. Cěch–de Rham formalism
[…]
$\hat{\delta}:\prod\Omega^{p}(U_{i_{0}\dots i_{m}})$ $\to\prod\Omega^{p}(U_{i_{0}\dots i_{m+1}})$
$(\hat{\delta}\omega)_{i_{0}\dots i_{m+1}}$ $\coloneqq\sum_{s=0}^{m+1}(-1)^{s}\,\omega_{i_{0}\dots\hat{i}_{s}\dots i_{m+1}},$
where the caret in the subscript means omission, and we only consider non-trivial intersections of orders $m$ and $m+1$ in the above definition. The rows in the diagram of Figure 8 are exact, except the last row
$\check{C}^{0}(\{U_{i}\},\mathbb{R})\xrightarrow{\quad}\check{C}^{1}(\{U_{i}\},\mathbb{R})\xrightarrow{\quad}\check{C}^{2}(\{U_{i}\},\mathbb{R})\xrightarrow{\quad}\dots,$
with $\check{C}^{m}(\{U_{i}\},\mathbb{R})$ denoting the space of locally constant functions on non-trivial $m$-intersections of $\{U_{i}\}$. This sequence is a differential complex, and its homology is called the Cěch cohomology of the cover $\{U_{i}\}_{i=1}^{k_{0}}$. See [BT82, Chapter 8] for a detailed description of the Čech-de Rham formalism.
We start with a $3$-form $\phi\in\mathcal{E}_{1+k_{3}}$, the choice of which would be described later in this proof. As its restrictions $\{\phi_{i}\}_{i=1}^{k_{0}}$ are exact on $\{U_{i}\}_{i=1}^{k_{0}}$, there exist unique coexact $2$-forms $\{\psi_{i}\}_{i=1}^{k_{0}}$ (with appropriate boundary conditions described in the Hodge-decomposition (2.1)) such that $d\psi_{i}=\phi_{i}$ for each $i\in\{1,\dots,k_{0}\}$. Since $\psi_{i}$ minimises the $L^{2}$-norm among all the primitives of $\phi_{i}$ on $U_{i}$, the variational characterisation (5.4) gives
$\mu^{{(3)}}_{{1,\operatorname{\mathrm{ex}}}}(U_{i})\leq\frac{\|\phi_{i}\|^{2}_{L^{2}\Omega^{3}(M)}}{\|\psi_{i}\|^{2}_{L^{2}\Omega^{2}(M)}},\qquad\forall\,i\in\{1,\dots,k_{0}\}.$ (5.5)
[…]
$d\omega_{ij}=d\psi_{j}-d\psi_{i}=\phi-\phi=0,$
and since $H^{2}(U_{ij},\mathbb{R})=0$ for all $U_{ij}$ by hypothesis, there exist unique coexact $1$-forms $\{\eta_{ij}\}$ (with appropriate boundary conditions) on $\{U_{ij}\}$ such that $\omega_{ij}=d\eta_{ij}$. Like before, we have from the variational characterisation (5.4) that
$\mu^{{(2)}}_{{1,\operatorname{\mathrm{ex}}}}(U_{ij})\leq\frac{\|\omega_{ij}\|^{2}_{L^{2}\Omega^{2}(M)}}{\|\eta_{ij}\|^{2}_{L^{2}\Omega^{1}(M)}},$ (5.6)
[…]
Again, since $H^{1}(U_{ijk},\mathbb{R})=0$ for all $U_{ijk}$, there exist unique coexact functions $\{\hat{\nu}_{ijk}\}$ on $\{U_{ijk}\}$ such that $\nu_{ijk}=d\hat{\nu}_{ijk}$, and the variational characterisation gives
$\mu^{{(1)}}_{{1,\operatorname{\mathrm{ex}}}}(U_{ijk})\leq\frac{\|\nu_{ijk}\|^{2}_{L^{2}\Omega^{1}(M)}}{\|\hat{\nu}_{ijk}\|^{2}_{L^{2}\Omega^{0}(M)}}.$ (5.7)
[…]
$\hat{\eta}_{ij}=\sum_{l=1}^{k_{0}}\rho_{l}\,\hat{\nu}_{lij}$
[…]
$\{\bar{\eta}_{ij}\}\coloneqq\{\eta_{ij}\}-d\{\hat{\eta}_{ij}\},$
whence we have
$d\{\bar{\eta}_{ij}\}=d\{\eta_{ij}\}-d^{2}\{\hat{\eta}_{ij}\}=d\{\eta_{ij}\}$
and
[…]
$=0.$
[…]
$\tau_{i}\coloneqq\sum_{j}\rho_{j}\,\bar{\eta}_{ij}.$
Finally, we define $\{\bar{\psi}_{i}\}\coloneqq\{\psi_{i}\}-d\{\tau_{i}\}$, which satisfy
$d\{\bar{\psi}_{i}\}=d\{\psi_{i}\}-d^{2}\{\tau_{i}\}=d\{\psi_{i}\}$
and
[…]
$=0.$
In other words, $\{\bar{\psi}_{i}\}$ agree on all the non-trivial intersections $\{U_{ij}\}$ and thus give rise to a global $2$-form on $M$, say $\bar{\psi}$.
We now proceed to derive the required lower bound for the quotient
$\frac{\|\phi\|^{2}_{L^{2}\Omega^{3}(M)}}{\|\bar{\psi}\|^{2}_{L^{2}\Omega^{2}(M)}},$
using the upper bounds for $\mu^{{(3)}}_{{1,\operatorname{\mathrm{ex}}}}(U_{i})$, $\mu^{{(2)}}_{{1,\operatorname{\mathrm{ex}}}}(U_{ij})$ and $\mu^{{(1)}}_{{1,\operatorname{\mathrm{ex}}}}(U_{ijk})$ obtained so far. Firstly,
$\bar{\psi}_{i}$ $=\psi_{i}-d\tau_{i}$
$=\psi_{i}-d\sum_{j=1}^{k_{0}}\rho_{j}\bar{\eta}_{ij}$
$=\psi_{i}-d\sum_{j=1}^{k_{0}}\rho_{j}(\eta_{ij}-d\hat{\eta}_{ij})$
$=\psi_{i}-d\sum_{j=1}^{k_{0}}\rho_{j}\left(\eta_{ij}-d\sum_{k=1}^{k_{0}}\rho_{k}\hat{\nu}_{kij}\right).$
We have from (5.5) that
$\left\|\psi_{i}\right\|^{2}\leq\frac{\left\|\phi_{i}\right\|^{2}}{\mu^{{(3)}}_{{1,\operatorname{\mathrm{ex}}}}(U_{i})}\leq\frac{\left\|\phi\right\|^{2}}{\mu^{{(3)}}_{{1,\operatorname{\mathrm{ex}}}}(U_{i})}.$
Also,
$\left\|d\sum_{j=1}^{k_{0}}\rho_{j}\eta_{ij}\right\|^{2}$ $=\left\|\sum_{j=1}^{k_{0}}d\rho_{j}\wedge\eta_{ij}\pm\sum_{j=1}^{k_{0}}\rho_{j}\,d\eta_{ij}\right\|^{2}$
$\leq 2k_{0}\left[\sum_{j=1}^{k_{0}}\left\|d\rho_{j}\wedge\eta_{ij}\right\|^{2}+\sum_{j=1}^{k_{0}}\left\|\rho_{j}\,d\eta_{ij}\right\|^{2}\right]$
$\leq 2k_{0}\left[\sum_{j=1}^{k_{0}}c_{\rho}\left\|\eta_{ij}\right\|^{2}+\sum_{j=1}^{k_{0}}\left\|d\eta_{ij}\right\|^{2}\right]$
[…]
$=2k_{0}\sum_{j=1}^{k_{0}}\left(\frac{c_{\rho}}{\mu^{{(2)}}_{{1,\operatorname{\mathrm{ex}}}}(U_{ij})}+1\right)\left\|\psi_{i}-\psi_{j}\right\|^{2}$
$\leq 4k_{0}\sum_{j=1}^{k_{0}}\left(\frac{c_{\rho}}{\mu^{{(2)}}_{{1,\operatorname{\mathrm{ex}}}}(U_{ij})}+1\right)\left(\left\|\psi_{i}\right\|^{2}+\left\|\psi_{j}\right\|^{2}\right),$
and
$\left\|\sum_{j=1}^{k_{0}}d\rho_{j}\wedge\left(\sum_{k=1}^{k_{0}}d(\rho_{k}\hat{\nu}_{kij})\right)\right\|^{2}$ $\leq k_{0}\sum_{j=1}^{k_{0}}\left\|d\rho_{j}\wedge\left(\sum_{k=1}^{k_{0}}d(\rho_{k}\hat{\nu}_{kij})\right)\right\|^{2}$
$\leq k_{0}c_{\rho}\sum_{j=1}^{k_{0}}\left\|\sum_{k=1}^{k_{0}}d(\rho_{k}\hat{\nu}_{kij})\right\|^{2}$
$\leq k_{0}^{2}c_{\rho}\sum_{j=1}^{k_{0}}\sum_{k=1}^{k_{0}}\left\|d(\rho_{k}\hat{\nu}_{kij})\right\|^{2},$
with
$\left\|d(\rho_{k}\hat{\nu}_{kij})\right\|^{2}$ $=\left\|d\rho_{k}\,\hat{\nu}_{kij}\pm\rho_{k}\,d\hat{\nu}_{kij}\right\|^{2}$
$\leq 2\left[c_{\rho}\left\|\hat{\nu}_{ijk}\right\|^{2}+\left\|d\hat{\nu}_{ijk}\right\|^{2}\right]$
[…]
$\leq 2\left(\frac{c_{\rho}}{\mu^{{(1)}}_{{1,\operatorname{\mathrm{ex}}}}(U_{kij})}+1\right)\left\|\eta_{ij}\pm\eta_{ik}\pm\eta_{jk}\right\|^{2}$
$\leq 6\left(\frac{c_{\rho}}{\mu^{{(1)}}_{{1,\operatorname{\mathrm{ex}}}}(U_{kij})}+1\right)\left(\left\|\eta_{ij}\right\|^{2}+\left\|\eta_{ik}\right\|^{2}+\left\|\eta_{jk}\right\|^{2}\right)$
$\leq 6\left(\frac{c_{\rho}}{\mu^{{(1)}}_{{1,\operatorname{\mathrm{ex}}}}(U_{kij})}+1\right)\left(\frac{\left\|\psi_{i}\right\|^{2}+\left\|\psi_{j}\right\|^{2}}{\mu^{{(2)}}_{{1,\operatorname{\mathrm{ex}}}}(U_{ij})}+\frac{\left\|\psi_{i}\right\|^{2}+\left\|\psi_{k}\right\|^{2}}{\mu^{{(2)}}_{{1,\operatorname{\mathrm{ex}}}}(U_{ik})}+\frac{\left\|\psi_{j}\right\|^{2}+\left\|\psi_{k}\right\|^{2}}{\mu^{{(2)}}_{{1,\operatorname{\mathrm{ex}}}}(U_{jk})}\right).$
Using all the above estimates, we get
$\left\|\bar{\psi}_{i}\right\|^{2}\leq 3\left\|\psi_{i}\right\|^{2}+12k_{0}\sum_{j=1}^{k_{0}}\left(\frac{c_{\rho}}{\mu^{{(2)}}_{{1,\operatorname{\mathrm{ex}}}}(U_{ij})}+1\right)\left(\left\|\psi_{i}\right\|^{2}+\left\|\psi_{j}\right\|^{2}\right)\\ +18k_{0}^{2}c_{\rho}\sum_{j=1}^{k_{0}}\sum_{k=1}^{k_{0}}\left(\frac{c_{\rho}}{\mu^{{(1)}}_{{1,\operatorname{\mathrm{ex}}}}(U_{kij})}+1\right)\left(\frac{\left\|\psi_{i}\right\|^{2}+\left\|\psi_{j}\right\|^{2}}{\mu^{{(2)}}_{{1,\operatorname{\mathrm{ex}}}}(U_{ij})}+\frac{\left\|\psi_{i}\right\|^{2}+\left\|\psi_{k}\right\|^{2}}{\mu^{{(2)}}_{{1,\operatorname{\mathrm{ex}}}}(U_{ik})}+\frac{\left\|\psi_{j}\right\|^{2}+\left\|\psi_{k}\right\|^{2}}{\mu^{{(2)}}_{{1,\operatorname{\mathrm{ex}}}}(U_{jk})}\right).$
Employing inequality (5.5) and taking the summation on both sides with $i$ in the range $\{1,\dots,k_{0}\}$, we get
$\frac{\left\|\bar{\psi}\right\|^{2}}{\left\|\phi\right\|^{2}}\leq\frac{\sum_{i=0}^{k_{0}}\left\|\bar{\psi}_{i}\right\|^{2}}{\left\|\phi\right\|^{2}}\leq 18k_{0}^{2}\aleph,$
where $\aleph$ is as defined in the statement of Lemma 5.6. Finally, the variational characterisation (5.4) implies that
$\mu^{{(3)}}_{{1+k_{3},\operatorname{\mathrm{ex}}}}(M)\geq\frac{\|\phi\|^{2}_{L^{2}\Omega^{3}(M)}}{\|\bar{\psi}\|^{2}_{L^{2}\Omega^{2}(M)}}\geq\frac{(18k_{0}^{2})^{-1}}{\aleph}.$
Acknowledgement
The authors are grateful to Bruno Colbois for sharing his valuable insights and feedback on the preprint. Tirumala Chakradhar is a PhD student at the University of Bristol and thanks his supervisor Asma Hassannezhad for her guidance and the numerous discussions. The authors would also like to thank Alessandro Savo for the helpful feedback and suggestions.